### v1

# Audio Personas: Augmenting Social Perception via Body-Anchored Audio Cues

Abstract.
We introduce Audio Personas, enabling users to ”decorate” themselves with body-anchored sounds in audio augmented reality. Like outfits, makeup, and fragrances, audio personas offer an alternative yet dynamic channel to augment face-to-face interactions. For instance, one can set their audio persona as rain sounds to reflect a bad mood, bee sounds to establish personal boundaries, or a playful ”woosh” sound to mimic passing by someone like a breeze. To instantiate the concept, we implemented a headphone-based prototype with multi-user tracking and audio streaming. Our formative study with designers revealed that audio personas were preferred in public and semi-public-private spaces for managing social impressions (e.g., personality) and signaling current states (e.g., emotions). Our preregistered in-lab study with 64 participants showed that audio personas influenced how participants formed impressions. Individuals with positive audio personas were rated as more socially attractive, more likable, and less threatening than those with negative audio personas.
1. Introduction
In digital environments, such as social media or virtual reality, users can transform how others perceive them by adjusting the characteristics of their digital avatars. With just a few clicks, one can modify attributes like hairstyle, facial features, and body type–or even embody completely different identities by switching between genders or adopting non-humanoid avatars (Turkle, 2011).
While people can easily modify their social images online, face-to-face interactions lack such flexibility. Physical adornments like makeup and clothing remain relatively static, limited in their ability convey changing social messages. As such, there is a growing trend toward integrating digital elements, traditionally limited to virtual interactions, into real-world settings. One prominent example is the use of visual filters, popularized through social media, are now adopted in mobile augmented reality (mobile AR). By anchoring visual effects such as face masks or virtual jewelries to one’s body, it alters how one is socially perceived by others (Häkkilä et al., 2017; Roinesalo et al., 2016).
Despite growing efforts to enhance face-to-face interactions with digital cues, most explorations have concentrated on augmenting one’s visual appearance. However, our perception of others is inherently multisensory, involving not just how one looks, but also other sensory inputs such as their voice, footstep sounds, and their scent (Fiore, 1993; Fuller et al., 2007). In this work, we explore the opportunities to enhance face-to-face interactions through the auditory channel.
We introduce the concept of audio personas111Here, a persona refers to the aspect of someone’s character that is presented to or perceived by others (Oxford English Dictionary, 2023; Goffman, 2023)., body-anchored audio cues to augment social perception in face-to-face interactions. Sound effects are commonly used in online games to provide immediate feedback on user actions and enhance engagement (Liljedahl, 2011; Lipscombc, 2013; Kenwright, 2020). However, it’s potential use face-to-face social contexts is largely unexplored. In fact, psychology research has shown that mere exposure to emotion-evoking sounds can influence how people perceive the social attractiveness and emotional states of others in photographs (May and Hamilton, 1980; Hanser et al., 2015; Logeswaran and Bhattacharya, 2009). With the rise of spatial audio, the ability to spatially anchor sounds to individuals—similar to how visual filters personalize appearance—creates a novel means of personalized expression through sounds.
Figure 1 visualizes a few examples of audio personas: one can choose to surround themselves with bee sounds to set personal boundaries, a ”woosh” sound to mimic passing by like a breeze for fun and playfulness, or keyboard-clicking sounds to signal that they are in a busy state. To instantiate this concept, we first built a headphone-based prototype that utilizes multi-person tracking and multi-channel audio streaming. The user anchors sounds to their body and projects them to others in audio augmented reality.
We conducted two user studies to explore the concept of audio personas from two perspectives: (1) when users design and project audio personas, i.e., senders, and (2) when users listen to the audio personas of others, i.e., receivers.
Study #1 involved 8 designers to gain an initial understanding of the social intentions and contexts for using audio personas. We identified intended use cases for audio personas in public and semi-public-private spaces as a way to manage social impressions (e.g., personality) and signal current states (e.g., emotions). We also elicited the creations of 48 audio persona designs and found that audio personas being body-anchored is a key design feature. Additionally, we identified diverse design patterns in localization, triggers, dynamics and content of audio personas.
Study #2 built upon the identified use cases, examining how audio personas influence social impression and perceived emotional states from the receiver’s perspective. In this preregistered in-lab study (n = 64), we focused on key design factors of audio personas (i.e., body-anchored, audio valence). Participants interacted with two research assistants, while hearing the research assistants’ audio personas or hearing sounds anchored to objects in the space. We found that participants were more likely to incorporate audio personas, body-anchored audio cues, into their impressions of others compared to the object-anchored audio. Moreover, participants perceived the person with a positive audio persona as more socially attractive, likable and less threatening than the person with a negative audio persona. These findings align with the theoretical grounding that social perception is a multisensory process (Livesley and Bromley, 1973; Workman and Johnson, 1991). The spatial alignment of audio with a person, as introduced by audio personas, enhances the likelihood of the sound being perceived as relevant cues of that individual in social interactions.
Our key contributions are:
We introduce audio personas, body-anchored audio cues to enhance social perception in face-to-face interactions. The concept opens up possibilities for engaging dynamic, multisensory cues to shape social perception.
Through a formative study with designers, we explored the social intentions and contexts for audio persona usage. This study also uncovered design possibilities for audio personas and identified key design patterns.
In a preregistered in-lab study (n = 64), we showed that audio personas can influence how one forms impressions of others. These impressions were shaped in alignment with the valence of the audio persona.
2. Related Work
Our work builds on prior research that uses digital media to enhance social perception, as well as ongoing efforts to integrate dynamic digital elements into face-to-face interactions. We extend previous research that employs auditory cues to support physical interaction, i.e., audio augmented reality (audio AR).
2.1. Social Perception in Digital Media
In everyday life, people choose clothing, apply makeup, and use perfume to influence how others perceive them. This process, often referred to as impression management (Leary and Kowalski, 1990; Goffman et al., 2002), plays a crucial role in shaping social interactions. By selecting their attire and accessories, individuals convey different aspects of their self-identity (Lennon and Davis, 1989; Branstetter, 1975; Miller, 1997; Workman and Johnson, 1991; Croijmans et al., 2021).
Just as people carefully manage their social impressions in face-to-face interactions, they also do so in the digital world. On social media apps like Instagram and Snapchat (Choi and Sung, 2018), individuals craft their online presences via selfies and profile photos (Segalin et al., 2017; Siibak, 2009). Using digital filters and editing tools, users can easily enhance and transform appearances by smoothing skin, changing hair color, and adding digital accessories or makeup (Siddiqui, 2021; Javornik et al., 2022; Hong et al., 2020). Similarly, these digital filters have also been adopted in videoconferencing to augment remote communications (Li and Lee, 2023; Leong et al., 2021).
In virtual worlds, such as those found in games and virtual reality (VR) environments, users can create and represent themselves as avatars. On platforms such as Second Life (sec, 2024), self-avatars serving as users’ digital personas within the virtual space might or might not align with their physical social identity (Martey and Consalvo, 2011; Schultze and Leahy, 2009). Game designers also map sound effects to player movements to indicate the outcomes of their actions (Liljedahl, 2011; Lipscombc, 2013; Kenwright, 2020). These sounds provide immediate feedback, such as the noise of a sword hitting an enemy or the sound of footsteps, to enhance immersion. In first-person games such as Half-Life (hal, 2024), the introduction of avatar sound effects was found to increase gamer experience (Nacke et al., 2010).
With the emergence of social VR platforms, such as Horizon (hor, 2024) and VR Chat (vrc, 2024), users are able to directly embody the digital avatars that represent them (Kilteni et al., 2012). These avatars can not only reflect their desired appearance (Kim et al., 2023) but can also transcend their physical presence (Yee and Bailenson, 2007; Bailenson et al., 2004). For instance, users can become giants (Abtahi et al., 2019), have novel body parts (Won et al., 2015) or even adopt non-humanoid avatars (Krekhov et al., 2019; Ahn et al., 2016). There is also a growing attention to how audio in VR, such as changes to the avatar’s voice (Choi et al., 2023; Kao et al., 2022; Byeon et al., 2023; Cheng et al., 2024) or the presence of footstep sounds (Kern and Ellermeier, 2020) impact immersive experience. Similarly, audio is gaining attention in remote collaboration contexts (Yang et al., 2020; Takayama and Nass, 2010). Researchers found that collaborating with avatars with self-similar voices led to higher ratings on the characters’ likability and believability (Guo et al., 2024).
2.2. Introducing Dynamic Digital Elements to Face-to-Face Social Interaction
While digital media provides extensive flexibility in managing online social perception, the physical adornments we use every day are often static and do not adapt to evolving social contexts. Thus, researchers explored to introduce dynamic digital characteristics into physical interactions.
Drawing inspiration from the filters used in social media and online meeting platforms (Rios et al., 2018), researchers explored the application of visual filters in everyday augmented reality settings (Chakrabarty, 2024). For example, by attaching AR markers onto clothing, users can see animated or interactive elements superimposed onto their physical attire (Häkkilä et al., 2017; Roinesalo et al., 2016; KP et al., 2023). Additionally, mobile AR technology enables the digital modification of the appearance of physical jewelry and clothing, allowing users to experiment with different colors, patterns, and designs in real-time (Rantala et al., 2018; Roinesalo et al., 2016; Mackey et al., 2020; Fuste and Schmandt, 2018).
From a different angle, researchers also explored embedding wearable displays onto physical clothing to make physical adornments interactive and adaptive. An early prototype, BubbleBadge (Falk and Björk, 1999), was designed as a coat badge and was able to provide various information about the wearer. Similarly, accessories such as necklaces (Buruk et al., 2021), handbags (Colley et al., 2016), and rings (Buruk et al., 2021) can also be embedded with LEDs and small displays to change appearances dynamically. Researchers also explored technical possibilities of making it battery-free (Dierk et al., 2018), interfacing directly with the skin (Wang et al., 2017), and integrating with existing textile printing methods (Howell et al., 2016; Kan et al., 2015; Devendorf et al., 2016).
While our perception of others involve both visual (e.g., facial expressions, body language, clothing) and non-visual cues (e.g., voice, motion sounds, scents) (Fiore, 1993; Fuller et al., 2007), prior works have primarily focused on augmenting visual appearance. Our work broadens the range of modalities used to enhance physical social interactions with digital cues.
2.3. Dynamic Auditory Cues to Influence Social Perception
In this work, we propose using audio cues, anchored to the body, to augment social perception in face-to-face interactions. This approach is inspired by previous research in audio augmented reality (Yang et al., 2022), which demonstrated that altering environmental and body-generated sounds can affect how individuals perceive both their surroundings and bodies.
The term ”soundscape” refers to an acoustic environment perceived by human ears (Pijanowski et al., 2011). In audio augmented reality, soundscapes were used to deliver location-based sonic experiences to enhance the understanding of new sites (Rozier, 2000; McGookin and Brewster, 2012), support wayfinding (McGookin et al., 2009) and facilitate interactions with everyday objects (Yang et al., 2019). Sikora et al. found that visitors exposed to an ancient medieval soundscape (e.g., horse-riding sounds) at an archaeological site showed higher arousal than those without the soundscape. (Sikora et al., 2018). By introducing audio effects such as delay and frequency changes to environmental sounds (Storek et al., 2015; Watanabe and Terada, 2020), researchers developed systems to ”warp” the perception of the surrounding space.
Our bodies generate sounds both within themselves (e.g., breathing, heartbeat) and through interactions with the environment (e.g., tapping sounds). Tajadura-Jiménez et al. introduced the concept of ”sonic self-avatar,” emphasizing that self-representations in mixed reality environments should incorporate body-centered sounds (Tajadura-Jiménez et al., 2018; Tajadura and Västfjäll, 2008). When these body-centered sounds are altered, they can influence users’ mental representations of their bodies (Tajadura-Jiménez et al., 2015b; D’Adamo et al., 2024; Tonetto et al., 2014), perception of objects (Furfaro et al., 2015), and even gaits and motion (Tajadura-Jiménez et al., 2015a). For example, amplifying the high frequencies of footstep sound in walking can create the self-perception of having a thinner body (Tajadura-Jiménez et al., 2015a). These audio-based sensory feedback devices were further explored in the context of physical activity and rehabilitation (Ley-Flores et al., 2024; Turmo Vidal et al., 2024; Gomez-Andres et al., 2020).
Building on previous research that showed modifying environmental and body-centered sounds can enhance perceptions of space and self, this work taps into the exploration of audio in social contexts. Previous psychology research has shown that emotion-evoking music can influence how others are perceived in photographs, affecting factors like social attractiveness (May and Hamilton, 1980; Hanser et al., 2015), perceived emotional states (Logeswaran and Bhattacharya, 2009; Ignacio et al., 2021; Carroll and Young, 2005) and even interpersonal distance (Tajadura-Jiménez et al., 2011). In computer-mediated communication, audio serves as a form of social signaling. Previous research explored music-sharing systems (Golan, 2018; Mueller and Karau, 2002; Gurion and Jacoby, 2013), sonic identification to aid in remembering others (Choi et al., 2011; Choi, 2017) and background auditory cues linked to people’s activities (Mynatt et al., 1998). However, none of the previous work has explored the use of body-anchored sounds to facilitate dynamic social expressions in face-to-face interactions, opening up a wide range of design opportunities beyond what has been explored so far.
3. Proof-of-concept System
To deliver the audio persona experience, we implemented a proof-of-concept system that supports: (1) spatial and motion tracking of each user and (2) multi-channel audio streaming. The setup enables multi-user interactions, allowing users to both project and listen to audio personas. Figure 2 shows the system setup.
3.1. Spatial and Motion Tracking
We used HTC VIVE base stations and trackers 2.0 (HTC, 2024) for spatial and motion tracking. Each user wears a VIVE tracker on their forehead to approximate body location and orientation. Optionally, additional trackers could be attached to body joints (e.g., feet, hands) for limb tracking. Each tracker records 6DOF tracking data, including position () and orientation (). Tracking data was processed in real-time in Unity using the inverse kinematics pipeline, Final IK(fin, 2024), to create a digital twin of the user’s physical movements, as shown in Figure 2b. The audio persona was rendered by anchoring the audio source to a specific body joint of the user avatar (e.g., head, feet) within Unity. This setup supports customizable audio persona experiences, regardless of the number of trackers used.
Tracking the user’s head alone offers a wide range of interaction possibilities. Given the head locations of each user (), the system enables proximity-based interactions by calculating the distances between head trackers (, ). The system can then trigger audio personas, such as a bee sound, when users are at a certain distance. Moreover, by computing the change in head position () and orientation over time, the system can derive the speed of the user’s movement (e.g., ). This allows for compound effects: for instance, users can project a ”woosh” sound when they move fast and approach another user.
With additional body trackers, more fine-grained, movement-based audio personas can be implemented. For example, a footstep-based audio persona (e.g., giant stomps) involves assigning virtual colliders to the avatar’s feet () and the virtual floor (), pre-calibrated to match the actual floor height in Unity. Collisions are detected when the vertical position of a foot (): , where is collision detection threshold.
3.2. Audio Rendering and Multi-channel Audio Streaming
We further used Unity as the audio engine for the system, which provides control over 3D sound dynamics (e.g., min/max distance and volume rolloff). For realistic spatial audio rendering, we used the Meta XR Spatializer (met, 2024).
To support multi-user audio persona experience, we implemented multi-channel audio streaming. This means that each user using the system is a sender, broadcasting their audio persona to others while simultaneously hearing everyone else’s audio personas. To accomplish this, each user avatar in Unity was attached with one (or more) audio sources and one audio listener.
Specifically, we extended Unity’s default capabilities by implementing support for multiple virtual listeners, going beyond the standard single-listener setup. We built on top of the plugin developed by Gael Vanhalst (Mul, 2024), which provided a foundation for this capability. Additionally, we integrated the Audio Stream pipeline (Aud, 2024), which supports PCM audio streaming, to further route the audio output to separate audio devices connected to the host computer. Each audio persona the user hears is spatialized based on the location of the person who set that specific audio persona. In the current setup, the maximum number of users the system can support is limited by the number of output devices the host computer can connect to.
4. Study #1: User Perception and Designs of Audio Personas
What do people want to express with audio personas, and how do they use audio persona to convey information? To explore user perception and the design possibilities of audio personas, we conducted a formative study.
We chose to study participants with prior professional experience in audio. This sample selection allows us to understand the potential contexts and intentions behind audio persona usage from a user perspective, as well as elicit designs with expert insights.
4.1. Methods
4.1.1. Participants
We recruited 8 participants with prior experience in producing/working with audio content from diverse backgrounds: audio engineering (), VR/AR design (), music () and immersive arts (). Each participant had at least 2 years of experience in producing audio content and received training from university-level programs. The participants were recruited through word-of-mouth.
4.1.2. Procedure
Participants were first introduced to the concept of audio personas. They put on the prototype system, as described in Sec. 3, and interacted with an experimenter to experience three examples of audio personas: (1) a bee sound when approaching the experimenter, (2) a ”woosh” sound when the experimenter walked past, and (3) exaggerated stomping sounds from the experimenter. These demos intended to provide inspiration, and participants were encouraged to think beyond these audio persona designs.
We then interviewed participants about the social contexts where they believe audio personas could be useful. We began with open-ended questions and then discussed audio persona usage in different types of social contexts—public (e.g., festivals, parks, streets, public transport), semi-public-private (e.g., workspace, educational institutions, community events), and private (e.g., home, private gathering) (Karaçor, 2016; Swapan et al., 2019). We also explored social messages suitable for audio personas to express. Again, we began with open-ended questions and discussed preferences for expressing specific types of messages using audio personas, including emotions, ideas and opinions, needs and desires, traits and personality, identity and belonging, and personal experiences (Green, 2007).
Next, we elicited the creation of audio personas through two design sessions. Participants were asked to imagine themselves as designers of an audio persona library and create six different audio persona designs. In both sessions, participants used worksheets provided in the supplemental materials. In the first session, participants described what each audio persona should ”sound” like, its social contexts, and intended message. In the second design session, they specified sound characteristics, including spatial locations, triggers, and dynamics. At the end of each session, participants reviewed and discussed their designs with the experimenters.
The study took approximately 75 - 90 minutes, and participants were compensated $35 for their time. The study was approved by the Institutional Review Board (IRB # 71992).
4.1.3. Analysis
We video-recorded the sessions, transcribed the data, and organized it based on the guiding interview questions, as described in the previous section and shown in Figure 3. Two experimenters participated in coding the transcriptions and designs, and we identified common themes through inductive thematic analysis (Braun and Clarke, 2006). The themes identified through thematic analysis are reflected in the headings of the following subsections.
4.2. Finding #1: Social Intentions of Using Audio Personas
Figure 4 shows participants’ preferences for using audio personas to express different types of social messages. All participants agreed that audio personas are well-suited for expressing ”personality & traits” and ”emotions”, followed by ”identity & belonging.” In contrast, only 2 out of 8 found audio personas suitable for expressing ”needs & desires” or ”personal experience,” with ”ideas & opinions” as the least favored.
Together with qualitative findings, we identified two major social intentions for audio persona usage: managing social impressions (e.g., ”personality & traits,” ”identity & belonging”) and signaling current states (e.g., ”emotions”). We also identified the limitations of audio personas in conveying more complex information (e.g., ”needs & desires,” ”ideas & opinions,” ”personal experiences”).
4.2.1. Impression management
When discussing their intentions of using audio personas, P2, P3, P4 and P5 drew parallels with the selection of clothing. Similar to clothing, audio personas offer ”a way to express themselves” (P5), enabling others to ”get that first impression based on something that the other person set” (P2). Even more, P3 elaborated that audio persona are able to convey messages difficult to express through clothing:
”Somebody’s audio persona would tell you another dimension of how they think …some of them might be kind of cheeky or humorous and you can’t necessarily or harder to convey with clothing. Unless you buy one of those T-shirts that have a joke on it; and those get really old really quickly.”
As a dynamic medium, audio personas could serve as ”one of their intros when they’re meeting people” (P1), like a ”business cards” (P6), showing ”what they like, maybe what they don’t like” (P4), and fostering ”affinities” (P6). Moreover, audio personas were viewed as a means for creative expression. P8 described:
”This is a nice way to kind of bring your personality to something more fictional, less rooted in reality …you can really customize what other’s perception of you is …I feel like your physical presence feels different because of the sound.”
In the designs created by participants, 21 out of 48 audio personas intended to manage impressions, focusing on influencing how others perceive users. A wide range of audio content with varying valence was used and Figure 5 shows a selection of them. For instance, P7 designed an audio persona that features auditory accessories with a jingling bell sound to reflect personal jewelry taste and personality, as shown in (Figure 5a). P5 chose to augment their footstep sounds with a trombone sound, showing the love of SpongeBob (Figure 5b). P4 designed an audio persona that emits a dog barking sound whenever they carry their dog in a pouch, signaling that they are a pet owner and reminding others of the dog’s presence (Figure 5e).
4.2.2. Signaling current states
Another social intention for using audio personas is to signal the users’ current states and inform others on how they prefer to be interacted with. Audio personas offer an ”implicit” channel (Weiser and Brown, 1996) to convey current states, without saying out loud and ”imposing too much on [others]” (P4). P1 described it as ”another social cue in terms of emotions” and P5 saw it similar to ”setting your status” commonly found in chat apps.
P7 described examples of how they might use audio personas to signal current states. Different valences of sounds were leveraged to signal different emotional states.
”If I’m not feeling super chatty and I don’t want to engage in conversation with them, the [audio persona] can be something softer or more melodic, less exciting, more just like a kind of monotone. Versus like, if I’m feeling really high energy, it can be a more high energy sound”, said P7.
Such indications of one’s current state could help others to ”empathize better with where one’s at” (P3) and ”help other people to know how to interact with you” (P8). For instance, P2 described:
”If someone’s like, kind of down or depressed, like, we hear the rainfall?… maybe we could like, approach them and be like, ‘Oh, are you okay?’”
21 out of 48 designs were created to signal current states, reflecting users’ emotions, needs, or intended interactions, and Figure 6 shows a selection of them. For instance, P3 used the Mission Impossible theme song to signal urgency in settings like a grocery store, prompting others to step aside or assist (Figure 6d). In contrast, P4 designed peaceful forest sounds to indicate serenity and a desire not to be disturbed during activities like meditation (Figure 6b). P6 created a bombastic drum and metal guitar persona to express their dislike for chores such as dishwashing (Figure 6c).
4.2.3. Limitations in delivering complex information
When discussing messages that audio personas might not be suitable to convey, participants referred to those involving complex information or narratives, such as opinions and personal experiences. P2 attributed the limitations to how audio is prone to ”perceptual bias” and P5 stated these information need ”more clarity”. For instance, when considering ideas and opinions, P8 described them as ”very nuanced”. ”While audio is nuanced, it is always subjective. And so I don’t think it would clearly convey whatever idea or opinion would want it to pitch or rather, it has the potential to be misinterpreted”, said P8. Similarly, P4 expressed that personal experience is ”too complex a subject almost to capture in an audio clip”:
”That’s like you have to imply something that is reminiscent of your childhood or reminiscent of a recent life experience and hope that people capture that implied message; but I don’t think it’ll be as effective as compared to some of the other options.”
P1 suggested a good audio persona should be ”something that is not too long, like something that’s iconic, and then you could identify quickly.”
4.3. Finding #2: Social Context for Using Audio Personas
Figure 7 shows the participants’ rankings of their preferred social context for using audio personas. Public spaces (e.g., festivals, parks, streets, public transport) and semi-public-private spaces (e.g., workspace, educational institutions, community events) were found to be most preferred. Specifically, 6 out of 8 participants rated public spaces as where audio personas will be most useful for, followed by semi-public-private spaces. 7 out of 8 participants found the private space to be the least useful for using audio personas. In terms of audio persona designs, 36 designs were created for semi-public-private environments, 14 for public spaces, and 5 for private settings. The total exceeds 48 because some designs are intended for multiple contexts.
4.3.1. Intended for public and semi-public-private spaces
In a public context, P6 described their experience of communications as:”you’re not actually able to verbalize your thoughts to as many people as you want as comfortably”. Thus, audio personas can provide a nice way to ”get first impressions of strangers” (P2), serve as ”a novel way to meet and greet” (P3) and ”communicate some sort of affinity” (P6). Thus, users can potentially ”make new connections” as audio personas become something that people can ”bond over” (P4). P3 echoed, ”In public contexts, audio personas could help you network with new people that are similar to you”.
In semi-public-private spaces, P4 described the social interactions as: ”there are still gonna be a lot of strangers, and potentially coworkers.” Similarly, P2 described those spaces as more ”familiar” and they would ”become more open to different [audio] personas”. Similar to its use in public spaces, audio personas can serve as ”an interesting icebreaker …to get to know somebody else.” (P8). The usage of audio personas in semi-public-private space also involve ”understanding the vibes of somebody a little bit better” (P3). They elaborated with an example: ”When somebody is really zoned into something, they don’t want to be bothered at all …I can imagine, audio personas …being used for things like personal space violation.”
4.3.2. Less useful in private space
When considering a private social context, participants found it to be the least suitable for using audio personas. P1 described private gatherings as being: ”already pretty small. [and people are] in the same space and talking with each other.” P3 added that in those contexts, people usually ”know each other well enough”. Thus, participants deemed people should be able to ”start up a conversation” (P2), ”just talk to them honestly” (P4) and ”be fully present engaging with the person” (P8). In these contexts, audio personas were described as ”kind of unnecessary complication and distraction” (P3), making them feel ”less personal” (P5) and ”more disconnected from the people around them for conversation” (P7). On the other hand, P8 and P6 also described scenarios where they might want to use audio personas in a private context. P8 said: ”If I want to use it in a private space, it’s more of a novelty.” P6 envisioned an asynchronous usage where users re-experience other’s audio personas after private gatherings.
4.4. Finding #3: Design Patterns of Audio Personas
Lastly, we analyzed design patterns in the sound characteristics of audio persona designs, specifically examining their localization, activation triggers, and dynamic properties.
4.4.1. Localizations
We traced the sketches of audio persona locations marked by participants on the design worksheets and overlaid them to identify design patterns, as shown in Figure 8. One key observation across all 48 designs was that designers consistently anchored the audio persona to the user’s body, aligning the sound’s spatial location with the person’s. There was variability in the specific body parts where the audio personas were attached. 25 out of 48 designs were anchored to the upper body of the users, such as to the head (n = 11), hands (n = 4), and back (n = 3). 15 designs were placed around the lower body of the users, predominantly on the feet (n = 11). In 12 of 48 designs, participants circled the whole body instead of specifying a center origin, indicating they intended the audio personas to be heard around the body without a precise origin. 8 designs placed audio personas on multiple body parts.
While most designs were intended for individual use, P6 and P8 suggested that multiple people could share a group audio persona when engaging in similar activities, with each person contributing a part. In these cases, the audio persona would be placed at the group’s location, rather than from a single individual.
4.4.2. Triggers
Regarding what triggers the start of audio personas (i.e., when audio personas become audible to others), proximity is the most commonly used, as shown in Figure 9. Out of 48 designs, 30 were set to trigger when the distance between two users reached a predefined threshold. While most designs were triggered when others approached, in one of P6’s designs, they chose to trigger the audio persona only when someone walked past the user, calling it ”exit music.” 7 designs used specific gestures to trigger actions, such as the ”OK” hand gesture (P1) or air drumming motions (P7). 2 designs relied on gaze, like eye contact, as the trigger. 8 designs were context-dependent, activating in specific situations or when users engaged in certain activities, such as in a lecture hall (P6) or when a dog was nearby (P4). Additionally, participants explored using voice, body orientation (e.g., the direction the user is facing), or manual activation by the user. 8 designs explored a combination of multiple triggers.
4.4.3. Dynamics
Most audio personas designs from Study # 1 are dynamic, with evolving sound characteristics. 11 out of 48 designs involve a complete change in audio content based on factors like proximity, gestures, or user preferences. In one of P1’s designs, the audio persona shares the user’s favorite music, which adapts to the user’s changing taste. Other changes in dynamics involve the volume (n = 9), length (n = 7), pitch (n = 3), and speed (n = 2) of audio. For instance, in P2’s design, rain sounds vary in intensity to reflect the mood, becoming lighter as the person feels better. In 10 out of the 48 designs, participants chose to keep the audio persona unchanged.
When multiple people are present, the blending of their audio personas become possible. For example, in P8’s design, a group shares an audio persona (e.g., forest or river sounds) during a group meditation to create a collective sense of calm. Each user retains their own sound and, as more people join, additional sounds, like bird sounds, are introduced.
4.5. Summary of Findings in Study #1
In Study #1, we explored social context and intended social messages for using audio personas by working with designer participants. We identified potential use cases in managing social impressions (e.g., personality & traits, identity & belonging) and signaling current states (e.g., emotions) but are limited in communicating complex information (e.g., ideas & opinions). Participants preferred using audio personas in public (e.g., street, transport) and semi-private settings (e.g., workspace, educational institutions) to form bonds with people they do not know well, but found it less useful in private spaces. Analysis of 48 participant-generated audio personas revealed that anchoring sounds to the body is a key design feature. Additionally, a wide range of audio content with varied valence was used, and there are diverse design possibilities for localization, triggers, and dynamics.
5. Study #2: How Audio Persona Impacts Social Perception
Study #1 uncovered the intended use of audio personas to manage social impressions and signal current states. To explore whether audio personas are effective in influencing these dimensions of social perception, we need to examine from the receiver’s perspective. Does listening to others’ audio personas change how one perceive others?
In Study #2, 64 participants were invited to the lab for an icebreaker with two research assistants, with or without hearing the research assistants’ audio personas. As Study #1 identified body-anchored as a key design feature for audio persona and found that audio valence conveys distinct messages, Study #2 was designed around these two factors to examine the effect of audio personas. We hypothesized that hearing sounds anchored to a person (i.e., audio persona) would facilitate impression formation with the considerations of sounds. Furthermore, we hypothesized that the valence of the sounds would shape the direction of impressions and the perceived emotional states of others.
5.1. Condition Design
A mixed-factorial study design was adopted in the study, with one between-group and one within-group factor. The supplemental video illustrates the auditory experience of each condition.
5.1.1. Between-group: audio anchoring.
Audio personas being body-anchored was a key design feature observed in Study #1. Therefore, to understand the downstream impact of audio personas, it’s essential to understand how this spatial anchoring plays a role.
Specifically, participants were divided into two groups. Half of the participants were assigned to the audio persona group (Figure 10a, 10b), in which participants heard sounds anchored to each research assistants. As the research assistant walked past participants (e.g., moving from left to right or vice versa), their audio persona shifted in locations corresponding to their movements.
In the control group, participants interacted with research assistants without hearing their audio personas. Instead, participants heard sounds anchored to objects (i.e., speakers) in the space. We chose to include sounds in the control group as previous research showed that the mere presence of audio can already influence social emotional perception and behavior (Tajadura-Jiménez et al., 2011; Logeswaran and Bhattacharya, 2009; Carroll and Young, 2005). This control design thus better examine the exact effect of ”body-anchored” audio. Anchoring sound to objects is also commonly used in audio augmented reality (McGookin et al., 2009; Yang et al., 2019; Sikora et al., 2018) and the auditory experience in this group is similar to hearing sounds played from a physical speaker. The research assistants’ movements would not affect the sound.
5.1.2. Within-group: audio valence.
In Study #1, participants leveraged sounds with varying valences in audio persona designs to tailor different contexts, aligning with prior research that showed audio valence can evoke distinct emotions in listeners (Zentner et al., 2008; Juslin and Västfjäll, 2008; Tajadura and Västfjäll, 2008). Therefore, Study #2 also examined how the emotional component of the sounds used in audio personas influences social perception. Specifically, participants heard different types of sounds when interacting with each of the two research assistants. In the audio persona group, one research assistant had a positive audio persona, and the other had a negative audio persona.
In the control group, one object had a positive-valence sound (positive control), and the other had a negative-valence sound (negative control). Participants were instructed to stand on the opposite side from where the research assistants entered, ensuring they heard one sound while observing one assistant and a different sound while observing the other (Figure 10). This design was also used for the audio persona group to ensure consistency.
Nested factor: audio content. As Study #1 highlighted the wide range of potential audio content that could be used in audio persona designs, Study #2 considered the generalizability of the results by incorporating diverse sounds. Specifically, we included two different sounds for each valence condition (positive and negative), resulting in a collection of two positive and two negative sounds. For each trial, one positive and one negative sound were used.
We chose audio content from established audio databases to ensure control over the valence of sounds. Specifically, we used IADS-2 databse (Bradley and Lang, 2007) and its expanded version (IADS-E) (Yang et al., 2018), following prior work (Viinikainen et al., 2012; Masullo et al., 2021). We focused on nature and animal sounds due to their universal familiarity, and sounds were selected based on valence ratings provided in the database. The positive-valence sounds selected for the study were brooks (valence: 7.28) and forest (valence: 6.91), and negative-valence sounds selected were buzzing (valence: 1.92) and hurricane (valence: 2.18).
5.1.3. Counterbalancing
We used a Latin Square design to counterbalance the audio content participants heard when each research assistant was present. This resulted in 4 audio options for one research assistant and 2 options for the other one (covering two valence levels). Furthermore, we counterbalanced the side of the room each research assistant entered, which determined the participant’s standing location as described in Sec. 5.1.2. Therefore, the sample size needed to be a multiple of (audio option for research assistant #1) (audio option for research assistant #2) (side of room entering) (condition groups) = to account for all combinations across both groups. Additionally, we randomized the order in which each research assistant entered and the side of the room where each speaker was located.
5.2. Other Study Design Considerations
5.2.1. Study instruction
We introduced the study as an exploration of audio augmented reality for social interaction. Participants were informed that they would interact with two researchers and might hear sounds people chose to anchor to themselves or sounds anchored to objects in the space. The purpose of these sounds (e.g., signaling personality or emotions) was not disclosed to avoid biasing participants’ reactions with preconceived expectations, in line with prior work on the psychological impact of multisensory cues (Zhao et al., 2023; Costa et al., 2016).
5.2.2. Task design
Participants and two research assistants took part in an activity called ”Something’s Different” (Krueger, 2009). We adapted this common icebreaker task to foster spontaneous interactions. In the activity, participants needed to identify a change (e.g., clothing or accessories) in each research assistant between two sessions. In each session, research assistants entered the room individually, walking back and forth in front of the participants. The changes to each research assistant were always on their name tags to remain consistency across participants: one switched from a yellow name tag to blue, and the other from a blue name tag to yellow.
5.2.3. Research assistants training
All participants interacted with the same two male, Asian research assistants, who wore identical clothing across study sessions to minimize the influence of attire on social perception. We chose to prescript the research assistants’ behavior in the activity to minimize confounding factors and isolate the specific impact of the audio personas on social perception. Both research assistants were trained to maintain neutral body language and facial expressions in the study, similar to (Tajadura-Jiménez et al., 2011). The research assistants were blinded about the conditions or sounds assigned in all sessions to ensure consistent behavior.
5.2.4. Familiarization phase
Before the icebreaker, participants walked around the study space to familiarize themselves. In the control group, participants heard sounds as they approached the speakers on each side of the room, as they would later during the task. In the audio persona group, they did not hear anything since no research assistants were present. Participants identified the presence of sounds and their sources before proceeding with the task.
5.3. Apparatus
Figure 11 shows the study setup. The study space was 7 meters (length) x 5 meters (width) and was tracked by four HTC VIVE base station 2.0 units. Participants and research assistants all wore a head-mounted HTC 2.0 VIVE tracker for 6DOF head tracking at 80Hz. Two standing locations were marked on the floor, 2.4 meters apart. At each standing location, participants were positioned 1.5 meters from the audio source anchored to the nearest object. To maintain social comfort (Hall, 1966), the study space was arranged so that the minimum distance between participants and research assistants passing by was also approximately 1.5 meters.
Both body-anchored and object-anchored sounds were configured with a linear roll-off, with a minimum distance of 0.5 meters and a maximum distance of 2 meters. A 2-meter trigger distance was chosen based on audio persona designs from Study #1. Specifically, the audio persona group began hearing sound from a distance of 2 meters, with the volume increasing as the research assistants moved closer to them. The control group participants consistently heard the audio anchored to objects, regardless of the location of the research assistants. Details about sound characteristics are provided in the supplemental materials. Participants wore Bose QuietComfort Ultra Headphones in aware mode, allowing them to hear ambient sounds from the environment.
5.4. Procedure
Upon arriving at the lab, participants completed a consent form. Then, the experimenter introduced both research assistants to the participants by name and calibrated everyone’s head trackers based on their height. The participants were then introduced to the icebreaker task ”Something’s Different”, where their goal was to identify the change in each of the research assistants between two observation sessions. Before starting the task, researchers asked participants to walk around the study space to become familiar with it. Additionally, researchers showed participants the designated standing locations for observing the research assistants, as described in Sec. 5.1.
The task included two observation sessions. In each session, research assistants took turns walking back and forth in front of the participants five times. Only one research assistant was present in the room at a time. After both sessions, participants were instructed to turn to the table behind them and use the provided pen and paper to write down the changes they observed in each research assistant. Meanwhile, both research assistants entered the room again, and each sat on a bench in the space, as shown in Figure 11.
After writing their answers, participants were instructed to walk to each research assistant and show them the results. The research assistant responded with a nod after reading. Once participants had shown their answers to both research assistants, they were asked to place their answer sheets back on the table and return to sit on one of the benches. The experimenter then entered to reveal the correct answer to the change and guided the participants to another room to complete a post-task questionnaire.
The study was approved by the Institutional Review Board (IRB # 71992). The full study took approximately 30 minutes, and all participants were compensated $15 for their time.
Construct | Question |
likability | I think this person is competent. |
I think this person is likable. | |
I think this person is trustworthy. | |
perceived threat | I think this person is dominant. |
I think this person is threatening. | |
interpersonal attraction | I would enjoy completing a task with the person. |
I would like to interact with the person again. | |
I would like this person. | |
I would get along with this person. | |
perceived positive emotions | I think the person feels alert. |
I think the person feels determined. | |
I think the person feels inspired. | |
I think the person feels attentive. | |
I think the person feels active | |
perceived negative emotions | I think the person feels hostile. |
I think the person feels ashamed. | |
I think the person feels upset. | |
I think the person feels afraid. | |
I think the person feels nervous. |
5.5. Measures
We measured social impression and perceived current states of others, two major uses of audio personas identified in Study #1. We collected both self-report and behavioral data. In this manuscript, we focus on discussing the self-report measures, and all behavioral measures details are provided in the supplemental materials.
To understand how audio personas shape how one forms impressions of others, we started with impression drawings and impression writings. Drawing is a common technique in psychology to measure social perception (Kamphaus and Pleiss, 1991; Cherney et al., 2006), and we used drawing to understand if participants formed impressions of others with explicit consideration of their audio persona. This was contrasted by the drawings from the control group, where the sounds existed but were not spatially associated with the person. In support of the drawings, participants also wrote down their impressions of each research assistant using at least five sentences, referred to as impression writings. Two researchers independently coded the drawings and writings to analyze if participants referenced sounds through verbal or visual representations and resolved discrepancies through discussions.
To capture the nuances of social impressions beyond the incorporation of sounds, we included constructs such as likability judgments, threat perception, and interpersonal attraction, inspired by prior work (Fauville et al., 2022). All constructs were rated on a Likert scale from 1 (not at all) to 5 (very much), shown in Table 1. Likability judgments included the dimensions of attractiveness, competence, extroversion, likability, trustworthiness (Todorov et al., 2013; Fauville et al., 2022). Threat perception was measured by two dimensions, threat and dominance (Todorov et al., 2013; Fauville et al., 2022). Interpersonal attraction was measured by combining both social attraction (e.g., “I would get along with this person”) and task attraction (e.g., “I would like to interact with the person again.”), resulting in a 4-item construct (Fauville et al., 2022; Davis and Perkowitz, 1979).
To examine whether audio personas influence how people interpret the current states of others, we focused on the perceived emotions of others. Emotions were among the top-rated preferred messages participants would like to express using audio personas, as shown in Study #1. Specifically, we used the I-PANAS-SF scale (Karim et al., 2011), which includes 5 positive (active, attentive, alert, determined, inspired) and 5 negative terms (hostile, ashamed, upset, afraid, nervous), shown in Table 1.
We assessed reliability using Cronbach’s alpha for constructs with more than two items and Spearman-Brown formula for two-item constructs. Initial validation showed that reliability for likability (r) and threat perception (r) were lower than 0.7. A Principal Component Analysis identified two likability items (attractiveness, extroversion) loading in a different direction, which were dropped; this raised reliability to r. For the two-item threat perception construct, individual item results were reported, due to the inability to improve reliability via factor analysis.
5.6. Hypotheses
We preregistered the hypotheses through Open Science Foundation222https://osf.io/dzhca and outlined the key hypotheses below. We rephrased the hypotheses from the language originally used in the preregistration to ensure clarity in this paper.
Audio personas are body-anchored audio cues, creating spatial alignment between the person and the sound. From a multisensory cue combination standpoint (Spence, 2013; Zaki, 2013), such spatial alignment increases the likelihood of associating the sound as relevant information of that person. Thus, when considering its potential impact on impression formation, we hypothesized that audio persona group are more likely to incorporate visualizations of audio into their impression drawings of others, compared to control group, who heard sounds anchored to the objects (H1).
The valence of audio can trigger different emotional responses (Zentner et al., 2008; Juslin and Västfjäll, 2008; Tajadura and Västfjäll, 2008). Thus, when looking into the valence of sound used in audio personas, we hypothesized that people would form a more positive impression of a person with a positive audio persona compared to a person with a negative audio persona (H2). Even though control group didn’t experience sound spatially anchored to the person, the temporal alignment introduced by hearing sounds while interacting with a person might already have an effect on social perception, based on prior work (Tajadura-Jiménez et al., 2011; Logeswaran and Bhattacharya, 2009; Carroll and Young, 2005). Thus, we also hypothesized that people would form a more positive impression of others when there is a positive sound during their interactions, compared to when there is a negative sound, with the sounds anchored to objects in the environment (H3).
When comparing social impressions and perceived states between the two groups, we evaluate whether the spatial binding introduced by audio personas could enhance the effect of audio valence. We hypothesized a significant interaction between the audio anchoring and audio valence variables. Specifically, we hypothesized that a person with a negative sound anchored to them (i.e., negative audio persona) would create a more negative impression on others compared to when the negative sound is anchored to an object (H4). Similarly, we hypothesized that a person with a positive sound anchored to them (i.e., positive audio persona) would create a more positive impression on others compared to when the positive sound is anchored to an object (H5).
5.7. Participants
Following preregistration, we recruited 66 participants who were over 18 years old and had no hearing deficit. Two participants’ data were excluded from the following analysis because of technical errors. Thus, the final sample size is 64 participants: 36 self-identified as females, 27 self-identified as males, and 1 self-identified as non-binary. The average age of participants was 24.7 years old (). 14 participants identified as Caucasian, 3 identified as African American, 2 identified as Native American, 43 identified as Asian, 7 chose ”other” and 2 chose ”prefer not to say”. The total added to more than the sample size as participants could select more than one ethnicity. None of the participants knew the research assistants before the study.
5.8. Finding #1: Inclusion of Audio in Impression Formation
We found that 20 out of 32 (62.5%) participants in audio persona group explicitly described the sounds they heard during interactions with others in their impression drawings or writings. Of these, 11 visualized the sounds in impression drawings, 18 mentioned them in impression writings, and 9 included descriptions of sounds in both.
In contrast, only 6 out of 32 (18.75%) participants in control group explicitly described the sounds they heard during interactions with others in their impression drawings or writings. Of these, 2 visualized the sounds in impression drawings, 6 mentioned them in impression writings, and 2 included descriptions of sounds in both.
We applied a Z-test to compare the proportions of participants that include sounds in their impression drawings or writings of others. We found that the audio persona group were significantly more likely to include visualizations of sounds in the impression drawings compared to the control group (). Similarly, audio persona group were significantly more likely to describe sounds in the impression writings than control group (). These findings confirm H1.
Figure 12 shows a selection of impressions drawings made by participants in both groups. In audio persona group, participants often visualized the content of the sounds they heard, such as ”forest” or ”water” associated with positive audio personas, and ”bees” or ”hurricane” for negative audio personas. In contrast, in the control group, when the sounds were anchored to objects, participants tended to focus on drawing the person’s visual appearance.
Word count of ”sound”, ”sounds” and ”audio” in impression writings | ||
control group | audio persona group | |
with positive sound | 4 | 23 |
with negative sound | 5 | 24 |
total | 9 | 47 |
As an exploratory measure, we analyzed the frequency of the words ”sound,” ”sounds,” or ”audio” in participants’ impression writings (details in Table 2). These terms appeared 47 times in the audio personas group but only 9 times in the writings from the control group. Aligning with the impression writings, participants were more likely to incorporate sounds into their impressions of others when sounds were anchored to the person (i.e., audio personas), compared to when sounds were anchored to objects.
Looking into the content of impression writings, the audio persona group tended to describe how the sounds influenced their impressions of others. For instance, P2 described the research assistant with a positive audio persona: ”I liked when he passed by me since the sound of the water gently splashing was nice and calming.” Similarly, P62 wrote ”His sounds were like nature, which were complemented by his brown jumper and green hat. Together, that is peaceful”. When recalling the person with a negative audio persona, P24 wrote:”He seems like a nice guy, although the buzzing sound wasn’t as pleasant. The sound, along with his more defined facial structure like chin and cheekbones made him seem like a sharper guy who makes more judgments.” P26 wrote:”The clanking droid-like sounds inspires images of bluntness, stone faces, and assertiveness.”
In the control group, the impression writings were much more focused on the visual appearance of the research assistants. ”He was also walking evenly and with purpose. Maybe he’s chill, I might’ve gotten that impression from his clothes? He also might be interested in sports or fitness because he had a sports band and water bottle”, wrote P12.
5.9. Finding #2: Impact on Social Impression
Figure 13 and Figure 14 present the results related to social impressions. We used mixed ANOVA with audio anchoring as a between-group factor and audio valence as a within-group factor. This analysis was applied to measures including ratings of social attraction, threat, and likability as well as sentiment intensity in impression writing. While the two sounds within the valence condition have slight variation across the dependent variables, the trend remained the same and we opted to follow our preregistered plan of averaging across them within the condition, for ease of interpretation.
5.9.1. Social attraction.
Figure 13 presents the ratings for social attraction. A main effect was found for audio valence (), with a large effect size (). No main effect was found for audio anchoring () and no significant interaction was observed ().
While the interaction effect was not statistically significant as predicted (H4, H5), a significant main effect was observed for the audio valence. To identify the specific differences within each condition group regarding audio valence levels, as outlined in our preregistered hypotheses, we conducted exploratory pairwise comparisons, following prior work (Kocur et al., 2021; Johnson et al., 2023; Kirchner-Krath et al., 2024). Using Bonferroni-corrected pairwise T-tests, we found a significant difference () in social attraction ratings within the audio persona group. Participants rated the person with a positive audio persona () as more socially attractive than one with a negative persona (), confirming H2.
In the control group, the trend aligned with what we found in the audio persona group, but it was not statistically significant (), failing to support H3. When sounds were anchored to objects, the person was rated slightly more socially attractive with a positive sound () than with a negative sound ().
5.9.2. Perceived threat.
On the threat dimension, a main effect was found for audio valence () with a medium effect size (). No main effect was found for audio anchoring (), and no interaction effect was observed (). As with social attraction ratings, we conducted exploratory pairwise comparisons to examine differences within main effects. Post-hoc pairwise comparisons showed a significant difference in threat perception ratings between valence levels in the audio persona group (), with participants rating the person with a negative audio persona () as more threatening than the person with a positive audio persona (). In the control group, a similar trend appeared, though it was not significant (). When sounds were anchored to objects, the person was judged as slightly more threatening with a negative sound () compared to the person with a positive sound ().
On the dominance dimension, a main effect was found for audio valence (), with a medium effect size (). No main effect was found for audio anchoring (), nor the interaction effect (). Post-hoc pairwise comparisons showed no significant difference in dominance perception ratings between valence levels in the audio persona group () or control group (). However, the trend still aligned with the threat perception ratings. In audio persona group, participants rated the person with a negative audio persona () as slightly more dominant than the person with a positive audio persona (). When sounds were anchored to objects, the person was judged as slightly more dominant with a negative sound () compared to the person with a positive sound ().
5.9.3. Likability judgments.
A main effect was found for audio valence () on perceived likability, with a medium effect size (). No main effect was found for audio anchoring (), and no interaction effect was observed (). Post-hoc pairwise comparisons showed a significant difference in likability perception ratings between valence levels in the audio persona group (), with participants rating the person with a positive audio persona () as more likeable than the person with a negative audio persona (). In the control group, the trend aligned with audio persona group, though it was not significant (). When sounds were anchored to objects, the person was judged as slightly more likable with a positive sound () compared to a negative sound ().
5.9.4. Sentiments in impression writing.
As an exploratory measure, we conducted sentiment analysis on impression writings. We chose a lexicon-based model (NLTK Vader sentiment analyzer (Hutto and Gilbert, 2014)), due to its emphasis on sentiment intensity, which provides a more granular and interpretable analysis compared to classification-based deep-learning models (such as (Barbieri et al., 2020)). Despite its simplicity, we found it aligned well with human judgment.
Sentiment intensity results are shown in Figure 14. We found a main effect for audio valence () on the sentiment intensity, with a medium effect size (). No main effect was found for the audio anchoring (, ) and no interaction effect was observed (, ).
We conducted exploratory post-hoc pairwise comparisons to reveal significance in different audio valence levels, as in the previous measures. We found that the sentiment was significantly more positive () in impression writings describing the person with a positive audio persona () compared to writings about the person with a negative audio persona (). Control group had the same trend, though it was not significant (). When sounds were anchored to objects, the sentiment was slightly more positive in impression writings describing the person with a positive sound () than those describing the person with a negative sound ().
5.10. Finding #3: Impact on Perceived Emotional States of Others
Figure 15 shows the results for perceived emotional states of others. The analysis revealed no significant main effects of either audio anchoring (F) or audio valence () on the perceived positive emotions in others. Similarly, no significant main effects was found for either audio anchoring () or audio valence () on the perceived negative emotions in others. No significant interaction was found between audio anchoring and audio valence for perceived positive emotions () or perceived negative emotions (). This indicates that participants generally evaluated the emotional states of both of their partners similarly, regardless of the audio valence or anchoring.
6. Discussion
In this section, we discuss insights from our studies, limitations of this paper, and future directions for audio personas.
6.1. Sound, Impression Management and Social Perception
In Study #1, we studied with designer-participants and identified potential use cases of audio personas to manage social impressions (e.g., personality), and signal current states (e.g., emotions). Study #2 examined the effectiveness of audio persona in influencing these dimensions from the receiver’s standpoint. Here, we discussed key insights from the findings of both studies.
Study #1 revealed that body anchoring is a key feature for audio personas, consistently present in all 48 designs created by designer-participants. Study #2 further showed that such spatial anchoring of sound shaped how social impressions were formed. Participants explicitly integrate audio personas into their impressions of others, revealed in the impression writings and drawings. This finding was contrasted with control group, where participants heard the same sounds anchored to objects and were much less likely to integrate sounds into their impressions of others. These findings align with a theoretical framework of impression formation (Livesley and Bromley, 1973; Workman and Johnson, 1991), which suggests that individuals form impressions by focusing on specific cues from the available information about a person and making inferences based on those cues. From a multisensory cue combination standpoint (Spence, 2013; Zaki, 2013), an audio persona provides both spatial and temporal alignment between an auditory cue and a person, strengthening the association of sound as available information about the person. In contrast, in the control group, sounds anchored to objects had only temporal alignment (i.e., the sound played when the person was present) but lacked spatial alignment with the person. Such a multisensory integration process could explain impression formation in the context of audio personas.
Study #1 also showed that participants leveraged different valences of sound to convey various messages. Study #2 also investigated the emotional component of audio persona. We found participants rated the person with a positive audio persona as more socially attractive, likable, and less threatening than the person with a negative audio persona. We observed a similar trend in the control group, where sound was anchored to objects during interactions with others, although the difference was not statistically significant. These results aligned with prior work that showed emotion-evoking music can influence social impressions of individuals in photographs in the same direction of the music valence (May and Hamilton, 1980; Hanser et al., 2015).
We did not observe a significant interaction between audio anchoring and audio valence as originally hypothesized. We explicitly chose not to use a control group with no sound in order to measure the clear effect of audio anchoring. The subtlety of the manipulation suggests that the mere presence and temporal binding of audio may elicit a similar level of affective response, regardless of where the sound is spatially anchored. However, body-anchored sound does enhance the conscious integration of sound into impressions of the person, as demonstrated in the impression writings and drawings. This distinction may become especially important in multi-person interaction contexts, where individuals engage with several people simultaneously, unlike the one-on-one interactions explored in this paper. In such settings, body-anchored sounds may make a clearer distinction, potentially leading to more differentiated social impressions of each person, compared to situations where multiple sounds are simultaneously present in the space.
While previous research suggested that sounds prime the perceived emotions of others (Logeswaran and Bhattacharya, 2009; Ignacio et al., 2021; Carroll and Young, 2005), our study did not observe a similar effect. One possible explanation is that our study put participants into more complex social settings (beyond observing static faces as in prior work), where interpreting perceived emotions of others involve processing other cues (Lange et al., 2022). Since most non-verbal cues were controlled and kept neutral in our study, participants likely perceived the primary signals as neutral, overriding the priming effect of audio. This indicates that semantic congruence (Laurienti et al., 2004; Kim and Lee, 2022) between audio personas and other contextual information might be necessary for influencing the perceived emotional states of others.
6.2. Limitations
This paper’s investigations serve as key initial steps toward exploring audio personas but also come with several limitations.
On the system side, the current prototype relies on base stations and optical trackers. While it was sufficient for in-lab investigations of the concept, the existing system restricts interactions to equipped spaces and users wearing them. A more portable, user-friendly form factor is needed for real-world scalability. Future work could explore ultra-wideband (UWB) or computer vision for spatial tracking and leverage mobile devices for audio output. These improvements would enhance the accessibility and practicality of audio persona systems in everyday settings.
Study #1 interviewed designer-participants to gain initial insights of audio personas usage, following prior work (Liu et al., 2023). A larger and more diverse sample, particularly participants without a technical background, is needed to understand the preferences of the broader population. Additionally, while Study #1 elicited an initial set of designs, it did not fully explore the challenges of scaling audio personas, such as in large conference settings. Future studies should invite designers to investigate techniques like audio mixing for effective deployment in large-scale, multi-user contexts.
Study #2 employed a controlled, in-lab experiment to investigate how audio personas influence the perception of others, exploring the potential uses identified in Study #1 from the receiver’s perspective. While this design provided valuable insights into how key design factors of audio personas (e.g., body-anchored, audio valence) influence social perception, it has a few limitations. First, the behavior of research assistants was pre-scripted, which ensured experimental control, but did not fully capture the nuance of everyday interactions. Future work should examine audio persona use in more socially complex settings. For instance, researchers can incorporate other communication channels (e.g., spontaneous conversations) to explore how audio personas influence impression formation with additional contextual cues. Similarly, researchers should investigate the effects of audio personas without the explicit framing of ”studying how audio impacts social interactions” to encourage naturalistic reactions. Second, the study used object-anchored sounds as the control condition, which allows us examine the exact effect of body-anchoring in audio personas. However, future work should compare audio persona with other baseline, such as no audio or non-spatialized sound. Third, in Study #2, we used only nature and animal sounds, but a wider variety of content should be explored, such as human and mechanical sounds. This study also only focused on a single style of sound characteristics, investigating variations in localization, triggers, and dynamics of audio personas will provide new insights of how audio personas influence social perception. Studying a broader range of audio persona designs will also help us to understand their differences in affecting social perception. Ultimately, the effects of audio personas in multi-user contexts and real-world social interactions beyond the lab should be explored.
6.3. Future Directions
To deploy audio personas in real-world, multi-user contexts, we lay out several future directions.
Social identities are multifaceted and adapt fluidly based on contexts (Tajfel, 1974), suggesting that users may wish to have multiple audio personas. This opens up opportunities to explore interaction techniques that enable users to seamlessly create, project, and switch between audio personas tailored to different contexts. Additionally, as users both project and receive audio personas, the control over auditory experiences becomes essential. This includes the ability to choose which audio personas to hear and to start or stop listening as desired. Future research could investigate user preferences and interface designs that offer such control, such as opt-in and opt-out mechanisms.
While audio personas were suggested to be useful in many contexts in Study #1, more investigations are needed to understand situations where audio personas should not be used. For instance, when the user needs to focus on something auditory and not be distracted by other signals, audio personas might not be well-suited. Additionally, quantitative understanding is needed to assess the effectiveness of audio personas in conveying various types of messages. This would help users leverage audio personas to efficiently communicate intended messages and avoid miscommunication that might be introduced by audio-only interactions.
As future work explores more complex audio personas, leveraging knowledge in psychoacoustics offers great potential. For example, the best human auditory localization performance occurs within - in the horizontal plane and - in elevation (Carlini et al., 2024). In Study #1, we observed that designers localized audio personas to specific body parts to create motion-specific experiences, aligning well with these perceptual thresholds. Further research could investigate computational models to support the creation of perceptually distinct audio personas. This becomes particularly important to prevent sensory confusion in multi-user contexts.
Lastly, we recognize that the value of audio personas depend on how widely they are adopted. We also acknowledge that the current headphone-based prototype poses a technology barrier, limiting the augmented social experience to users with headphones. However, audio personas do not have to be a headphone-delivered experience. For example, personal speaker-based (e.g., ”boomboxes,” (Boyer, 2014) popular in the 1980s and 90s) or clothing-integrated audio personas could achieve more universally accessible experiences. We see diverse opportunities to deploy audio personas and examine their implications for social interactions.
7. Conclusion
We introduced Audio Personas, a novel concept to augment face-to-face interactions with body-anchored audio cues. Through a formative study with designers, we uncovered potential use cases and elicited design patterns. Our in-lab study further revealed that audio personas can influence how people formed social impressions. Individuals with a positive audio persona were deemed as more socially attractive, likable, and less threatening than ones with a negative audio persona. Audio personas open new possibilities for engaging dynamic multisensory channels to enhance face-to-face social interactions and this paper laid out key building blocks for future explorations.
Acknowledgements.
We would like to thank Keshav Rastogi, Muhammad Ali, and Brian Beams for their support on the user study logistics. We thank Anna Matsumoto and Emmanuel Corona-Moreno for their help on the video production. We would also like to thank everyone who provided feedback on the paper draft: Danyang Fan, Han Guo, Eugy Han, Saehui Hwang, Monique Santoso, Tara Srirangarajan, Olivia Tomassetti, Ahad Rauf, Savannah Cofer, Portia Wang, Yifei Cheng, Yiran Zhao, and Ran Zhou. The lead author of this work is supported by the Knight-Hennessy Scholarship.References
- (1)
- Aud (2024) 2024. Audio Stream. https://assetstore.unity.com/packages/tools/audio/audiostream-pcm-audio-for-unity-65411.
- hal (2024) 2024. Half-Life. https://www.half-life.com/en/home/.
- HTC (2024) 2024. HTC VIVE. https://www.vive.com/.
- hor (2024) 2024. Meta Horizon. https://horizon.meta.com.
- met (2024) 2024. Meta XR Audio SDK. https://assetstore.unity.com/packages/tools/integration/meta-xr-audio-sdk-264557.
- Mul (2024) 2024. Multi Audio Listener. https://github.com/GaelVanhalst/MultiAudioListener.
- fin (2024) 2024. RootMotion Final IK. https://assetstore.unity.com/packages/tools/animation/final-ik-14290.
- sec (2024) 2024. Second Life. https://secondlife.com.
- vrc (2024) 2024. VR Chat. https://hello.vrchat.com.
- Abtahi et al. (2019) Parastoo Abtahi, Mar Gonzalez-Franco, Eyal Ofek, and Anthony Steed. 2019. I’m a giant: Walking in large virtual environments at high speed gains. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. 1–13.
- Ahn et al. (2016) Sun Joo Ahn, Joshua Bostick, Elise Ogle, Kristine L Nowak, Kara T McGillicuddy, and Jeremy N Bailenson. 2016. Experiencing nature: Embodying animals in immersive virtual environments increases inclusion of nature in self and involvement with nature. Journal of Computer-Mediated Communication 21, 6 (2016), 399–419.
- Bailenson et al. (2004) Jeremy N Bailenson, Andrew C Beall, Jack Loomis, Jim Blascovich, and Matthew Turk. 2004. Transformed social interaction: Decoupling representation from behavior and form in collaborative virtual environments. Presence: Teleoperators & Virtual Environments 13, 4 (2004), 428–441.
- Barbieri et al. (2020) Francesco Barbieri, Jose Camacho-Collados, Luis Espinosa Anke, and Leonardo Neves. 2020. TweetEval: Unified Benchmark and Comparative Evaluation for Tweet Classification. In Findings of the Association for Computational Linguistics: EMNLP 2020. Association for Computational Linguistics, Online, 1644–1650. https://doi.org/10.18653/v1/2020.findings-emnlp.148
- Boyer (2014) BiLL Bahng Boyer. 2014. Urban echoes: The boombox and sonic mobility in the 1980s. The Oxford Handbook of Mobile Music Studies, Volume 1 1 (2014), 399.
- Bradley and Lang (2007) MM Bradley and PJ Lang. 2007. The international affective digitized sounds: affective ratings of sounds and instruction manual (Technical Report No. B-3). University of Florida. NIMH Center for the Study of Emotion and Attention, Gainesville, FL 30 (2007).
- Branstetter (1975) Katherine B Branstetter. 1975. ” I See by Your Outfit That You Are a Cowboy”: Attire as a Marker of Ethnic, Social, and Personal Identity. (1975).
- Braun and Clarke (2006) Virginia Braun and Victoria Clarke. 2006. Using thematic analysis in psychology. Qualitative research in psychology 3, 2 (2006), 77–101.
- Buruk et al. (2021) Oğuz’Oz’ Buruk, Çağlar Genç, İhsan Ozan Yıldırım, Mehmet Cengiz Onbaşlı, and Oğuzhan Özcan. 2021. Snowflakes: a prototyping tool for computational jewelry. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems. 1–15.
- Byeon et al. (2023) Hyeon Jeong Byeon, Seungjin Ha, and Uran Oh. 2023. Avocus: A voice customization system for online personas. In Extended Abstracts of the 2023 CHI Conference on Human Factors in Computing Systems. 1–6.
- Carlini et al. (2024) Alessandro Carlini, Camille Bordeau, and Maxime Ambard. 2024. Auditory localization: a comprehensive practical review. Frontiers in Psychology 15 (2024), 1408073.
- Carroll and Young (2005) Naomi C Carroll and Andrew W Young. 2005. Priming of emotion recognition. The Quarterly Journal of Experimental Psychology Section A 58, 7 (2005), 1173–1197.
- Chakrabarty (2024) Biswapriyo Chakrabarty. 2024. Exploring augmented reality solutions to facilitate co-located engagement. Ph. D. Dissertation. University of British Columbia.
- Cheng et al. (2024) Yi Fei Cheng, Laurie Heller, Stacey Cho, and David Lindlbauer. 2024. First or Third-Person Hearing? A Controlled Evaluation of Auditory Perspective on Embodiment and Sound Localization Performance (ISMAR ’24). IEEE.
- Cherney et al. (2006) Isabelle D Cherney, Clair S Seiwert, Tara M Dickey, and Judith D Flichtbeil. 2006. Children’s drawings: A mirror to their minds. Educational psychology 26, 1 (2006), 127–142.
- Choi et al. (2023) Minsoo Choi, Alexandros Koilias, Matias Volonte, Dominic Kao, and Christos Mousas. 2023. Exploring the Appearance and Voice Mismatch of Virtual Characters. In 2023 IEEE International Symposium on Mixed and Augmented Reality Adjunct (ISMAR-Adjunct). IEEE, 555–560.
- Choi and Sung (2018) Tae Rang Choi and Yongjun Sung. 2018. Instagram versus Snapchat: Self-expression and privacy concern on social media. Telematics and informatics 35, 8 (2018), 2289–2298.
- Choi (2017) Yongsoon Choi. 2017. Designing Wearable Electronic Spectacle for Enhancing Social Impressions Emitting Sound and Smell. International Journal of Multimedia and Ubiquitous Engineering 12, 4 (2017), 95–108.
- Choi et al. (2011) Yongsoon Choi, Adrian David Cheok, Xavier Roman, The Anh Nguyen, Kenichi Sugimoto, and Veronica Halupka. 2011. Sound perfume: designing a wearable sound and fragrance media for face-to-face interpersonal interaction. In Proceedings of the 8th international conference on advances in computer entertainment technology. 1–8.
- Colley et al. (2016) Ashley Colley, Minna Pakanen, Saara Koskinen, Kirsi Mikkonen, and Jonna Häkkilä. 2016. Smart handbag as a wearable public display-exploring concepts and user perceptions. In Proceedings of the 7th Augmented Human International Conference 2016. 1–8.
- Costa et al. (2016) Jean Costa, Alexander T Adams, Malte F Jung, François Guimbretière, and Tanzeem Choudhury. 2016. EmotionCheck: leveraging bodily signals and false feedback to regulate our emotions. In Proceedings of the 2016 ACM international joint conference on pervasive and ubiquitous computing. 758–769.
- Croijmans et al. (2021) Ilja Croijmans, Daniel Beetsma, Henk Aarts, Ilse Gortemaker, and Monique Smeets. 2021. The role of fragrance and self-esteem in perception of body odors and impressions of others. PloS one 16, 11 (2021), e0258773.
- D’Adamo et al. (2024) Amar D’Adamo, Marte Roel Lesur, Laia Turmo Vidal, Mohammad Mahdi Dehshibi, Daniel De La Prida, Joaquín R Diaz-Durán, Luis Antonio Azpicueta-Ruiz, Aleksander Väljamäe, and Ana Tajadura-Jiménez. 2024. SoniWeight Shoes: Investigating Effects and Personalization of a Wearable Sound Device for Altering Body Perception and Behavior. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–20.
- Davis and Perkowitz (1979) Deborah Davis and William T Perkowitz. 1979. Consequences of responsiveness in dyadic interaction: Effects of probability of response and proportion of content-related responses on interpersonal attraction. Journal of Personality and Social Psychology 37, 4 (1979), 534.
- Devendorf et al. (2016) Laura Devendorf, Joanne Lo, Noura Howell, Jung Lin Lee, Nan-Wei Gong, M Emre Karagozler, Shiho Fukuhara, Ivan Poupyrev, Eric Paulos, and Kimiko Ryokai. 2016. ” I don’t Want to Wear a Screen” Probing Perceptions of and Possibilities for Dynamic Displays on Clothing. In Proceedings of the 2016 CHI Conference on Human Factors in Computing Systems. 6028–6039.
- Dierk et al. (2018) Christine Dierk, Molly Jane Pearce Nicholas, and Eric Paulos. 2018. AlterWear: battery-free wearable displays for opportunistic interactions. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems. 1–11.
- Falk and Björk (1999) Jennica Falk and Staffan Björk. 1999. The BubbleBadge: a wearable public display. In Chi’99 extended abstracts on human factors in computing systems. 318–319.
- Fauville et al. (2022) Géraldine Fauville, Anna CM Queiroz, Mufan Luo, Jeffrey Hancock, and Jeremy N Bailenson. 2022. Impression formation from video conference screenshots: The role of gaze, camera distance, and angle. (2022).
- Fiore (1993) Ann Marie Fiore. 1993. Multisensory integration of visual, tactile, and olfactory aesthetic cues of appearance. Clothing and Textiles Research Journal 11, 2 (1993), 45–52.
- Fuller et al. (2007) Jerry Bryan Fuller, Tim Barnett, Kim Hester, Clint Relyea, and Len Frey. 2007. An exploratory examination of voice behavior from an impression management perspective. Journal of Managerial Issues (2007), 134–151.
- Furfaro et al. (2015) Enrico Furfaro, Frederic Bevilacqua, Nadia Berthouze, and Ana Tajadura-Jimenez. 2015. Sonification of virtual and real surface tapping: evaluation of behavior changes, surface perception and emotional indices. IEEE MultiMedia (2015).
- Fuste and Schmandt (2018) Anna Fuste and Chris Schmandt. 2018. ARTextiles for Promoting Social Interactions Around Personal Interests. In Extended Abstracts of the 2018 CHI Conference on Human Factors in Computing Systems (CHI ’18). ACM. https://doi.org/10.1145/3170427.3188589
- Goffman (2023) Erving Goffman. 2023. The presentation of self in everyday life. In Social theory re-wired. Routledge, 450–459.
- Goffman et al. (2002) Erving Goffman et al. 2002. The presentation of self in everyday life. 1959. Garden City, NY 259 (2002).
- Golan (2018) Amos Golan. 2018. IceBreakware: designing wearable technologies for spatial awareness and social interactions. Ph. D. Dissertation. Massachusetts Institute of Technology.
- Gomez-Andres et al. (2020) Alba Gomez-Andres, Jennifer Grau-Sánchez, Esther Duarte, Antoni Rodriguez-Fornells, and Ana Tajadura-Jiménez. 2020. Enriching footsteps sounds in gait rehabilitation in chronic stroke patients: a pilot study. Annals of the New York Academy of Sciences 1467, 1 (2020), 48–59.
- Green (2007) Mitchell S Green. 2007. Self-expression. OUP Oxford.
- Guo et al. (2024) Siqi Guo, Minsoo Choi, Dominic Kao, and Christos Mousas. 2024. Collaborating with my Doppelgänger: The Effects of Self-similar Appearance and Voice of a Virtual Character during a Jigsaw Puzzle Co-solving Task. Proceedings of the ACM on Computer Graphics and Interactive Techniques 7, 1 (2024), 1–23.
- Gurion and Jacoby (2013) Tom Gurion and Nori Jacoby. 2013. Audio-only augmented reality system for social interaction. In HCI International 2013-Posters’ Extended Abstracts: International Conference, HCI International 2013, Las Vegas, NV, USA, July 21-26, 2013, Proceedings, Part I 15. Springer, 322–326.
- Häkkilä et al. (2017) Jonna Häkkilä, Ashley Colley, Paula Roinesalo, and Jani Väyrynen. 2017. Clothing integrated augmented reality markers. In Proceedings of the 16th International Conference on Mobile and Ubiquitous Multimedia. 113–121.
- Hall (1966) Edward T Hall. 1966. The hidden dimension. Garden City (1966).
- Hanser et al. (2015) Waldie E Hanser, Ruth E Mark, Wobbe P Zijlstra, and Ad JJM Vingerhoets. 2015. The effects of background music on the evaluation of crying faces. Psychology of Music 43, 1 (2015), 75–85.
- Hong et al. (2020) Seoyeon Hong, Mi R Jahng, Namyeon Lee, and Kevin R Wise. 2020. Do you filter who you are?: Excessive self-presentation, social cues, and user evaluations of Instagram selfies. Computers in Human Behavior 104 (2020), 106159.
- Howell et al. (2016) Noura Howell, Laura Devendorf, Rundong Tian, Tomás Vega Galvez, Nan-Wei Gong, Ivan Poupyrev, Eric Paulos, and Kimiko Ryokai. 2016. Biosignals as social cues: Ambiguity and emotional interpretation in social displays of skin conductance. In Proceedings of the 2016 ACM Conference on Designing Interactive Systems. 865–870.
- Hutto and Gilbert (2014) Clayton Hutto and Eric Gilbert. 2014. Vader: A parsimonious rule-based model for sentiment analysis of social media text. In Proceedings of the international AAAI conference on web and social media, Vol. 8. 216–225.
- Ignacio et al. (2021) Daniel Andre Ignacio, David R Gerkens, Erick Ryan Aguinaldo, Dina Arch, and Ruben Barajas. 2021. Music-like stimuli affect the emotional processing of words and facial expressions. Psychology of Music 49, 3 (2021), 479–497.
- Javornik et al. (2022) Ana Javornik, Ben Marder, Jennifer Brannon Barhorst, Graeme McLean, Yvonne Rogers, Paul Marshall, and Luk Warlop. 2022. ‘What lies behind the filter?’Uncovering the motivations for using augmented reality (AR) face filters on social media and their effect on well-being. Computers in Human Behavior 128 (2022), 107126.
- Johnson et al. (2023) Janet G Johnson, Tommy Sharkey, Iramuali Cynthia Butarbutar, Danica Xiong, Ruijie Huang, Lauren Sy, and Nadir Weibel. 2023. UnMapped: Leveraging Experts’ Situated Experiences to Ease Remote Guidance in Collaborative Mixed Reality. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–20.
- Juslin and Västfjäll (2008) Patrik N Juslin and Daniel Västfjäll. 2008. Emotional responses to music: The need to consider underlying mechanisms. Behavioral and brain sciences 31, 5 (2008), 559–575.
- Kamphaus and Pleiss (1991) Randy W Kamphaus and Karen L Pleiss. 1991. Draw-a-Person techniques: Tests in search of a construct. Journal of School Psychology 29, 4 (1991), 395–401.
- Kan et al. (2015) Viirj Kan, Katsuya Fujii, Judith Amores, Chang Long Zhu Jin, Pattie Maes, and Hiroshi Ishii. 2015. Social textiles: Social affordances and icebreaking interactions through wearable social messaging. In Proceedings of the Ninth International Conference on Tangible, Embedded, and Embodied Interaction. 619–624.
- Kao et al. (2022) Dominic Kao, Rabindra Ratan, Christos Mousas, Amogh Joshi, and Edward F Melcer. 2022. Audio matters too: How audial avatar customization enhances visual avatar customization. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems. 1–27.
- Karaçor (2016) Elif Kutay Karaçor. 2016. Public vs. private: The evaluation of different space types in terms of publicness dimension. European Journal of Sustainable Development 5, 3 (2016), 51–51.
- Karim et al. (2011) Jahanvash Karim, Robert Weisz, and Shafiq Ur Rehman. 2011. International positive and negative affect schedule short-form (I-PANAS-SF): Testing for factorial invariance across cultures. Procedia-Social and Behavioral Sciences 15 (2011), 2016–2022.
- Kenwright (2020) Benjamin Kenwright. 2020. There’s more to sound than meets the ear: Sound in interactive environments. IEEE Computer Graphics and Applications 40, 4 (2020), 62–70.
- Kern and Ellermeier (2020) Angelika C Kern and Wolfgang Ellermeier. 2020. Audio in VR: Effects of a soundscape and movement-triggered step sounds on presence. Frontiers in Robotics and AI 7 (2020), 20.
- Kilteni et al. (2012) Konstantina Kilteni, Raphaela Groten, and Mel Slater. 2012. The sense of embodiment in virtual reality. Presence: Teleoperators and Virtual Environments 21, 4 (2012), 373–387.
- Kim and Lee (2022) Hayeon Kim and In-Kwon Lee. 2022. Studying the effects of congruence of auditory and visual stimuli on virtual reality experiences. IEEE Transactions on Visualization and Computer Graphics 28, 5 (2022), 2080–2090.
- Kim et al. (2023) Hayeon Kim, Jinhyung Park, and In-Kwon Lee. 2023. ” To Be or Not to Be Me?”: Exploration of Self-Similar Effects ofAvatars on Social Virtual Reality Experiences. IEEE Transactions on Visualization and Computer Graphics (2023).
- Kirchner-Krath et al. (2024) Jeanine Kirchner-Krath, Manuel Schmidt-Kraepelin, Sofia Schöbel, Mathias Ullrich, Ali Sunyaev, and Harald FO Von Korflesch. 2024. Outplay Your Weaker Self: A Mixed-Methods Study on Gamification to Overcome Procrastination in Academia. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–19.
- Kocur et al. (2021) Martin Kocur, Florian Habler, Valentin Schwind, Paweł W Woźniak, Christian Wolff, and Niels Henze. 2021. Physiological and perceptual responses to athletic avatars while cycling in virtual reality. In Proceedings of the 2021 CHI conference on human factors in computing systems. 1–18.
- KP et al. (2023) Nikita Menon KP, Eldy S Lazaro Vasquez, Hannah Curran, Sasha De Koninck, and Laura Devendorf. 2023. Weaving Augmented Reality Markers. In Adjunct Proceedings of the 2023 ACM International Joint Conference on Pervasive and Ubiquitous Computing & the 2023 ACM International Symposium on Wearable Computing. 310–314.
- Krekhov et al. (2019) Andrey Krekhov, Sebastian Cmentowski, and Jens Krüger. 2019. The illusion of animal body ownership and its potential for virtual reality games. In 2019 IEEE Conference on Games (CoG). IEEE, 1–8.
- Krueger (2009) Bonnie J Krueger. 2009. Ice breakers: Short activities to encourage social interaction. Lorenz Educational Press.
- Lange et al. (2022) Jens Lange, Marc W Heerdink, and Gerben A Van Kleef. 2022. Reading emotions, reading people: Emotion perception and inferences drawn from perceived emotions. Current opinion in psychology 43 (2022), 85–90.
- Laurienti et al. (2004) Paul J Laurienti, Robert A Kraft, Joseph A Maldjian, Jonathan H Burdette, and Mark T Wallace. 2004. Semantic congruence is a critical factor in multisensory behavioral performance. Experimental brain research 158 (2004), 405–414.
- Leary and Kowalski (1990) Mark R Leary and Robin M Kowalski. 1990. Impression management: A literature review and two-component model. Psychological bulletin 107, 1 (1990), 34.
- Lennon and Davis (1989) Sharron J Lennon and Leslie L Davis. 1989. Clothing and human behavior from a social cognitive framework Part I: Theoretical perspectives. Clothing and Textiles Research Journal 7, 4 (1989), 41–48.
- Leong et al. (2021) Joanne Leong, Pat Pataranutaporn, Yaoli Mao, Florian Perteneder, Ehsan Hoque, Janet M Baker, and Pattie Maes. 2021. Exploring the use of real-time camera filters on embodiment and creativity. In extended abstracts of the 2021 CHI conference on Human Factors in Computing Systems. 1–7.
- Ley-Flores et al. (2024) Judith Ley-Flores, Laia Turmo Vidal, Elena Márquez Segura, Aneesha Singh, Frederic Bevilacqua, Francisco Cuadrado, Joaquín Roberto Díaz Durán, Omar Valdiviezo-Hernández, Milagrosa Sánchez-Martin, and Ana Tajadura-Jiménez. 2024. Co-Designing Sensory Feedback for Wearables to Support Physical Activity through Body Sensations. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies 8, 1 (2024), 1–31.
- Li and Lee (2023) Benjamin J Li and Hui Min Lee. 2023. Filters uncovered: Investigating the impact of AR face filters and self-view on videoconference fatigue and affect. Telematics and Informatics Reports 11 (2023), 100088.
- Liljedahl (2011) Mats Liljedahl. 2011. Sound for fantasy and freedom. In Game sound technology and player interaction: Concepts and developments. IGI Global, 22–43.
- Lipscombc (2013) Scott D Lipscombc. 2013. Playing with sound: The role of music and sound effects in gaming. The Psychology of Music in Multimedia (2013), 289.
- Liu et al. (2023) Xingyu” Bruce” Liu, Vladimir Kirilyuk, Xiuxiu Yuan, Alex Olwal, Peggy Chi, Xiang” Anthony” Chen, and Ruofei Du. 2023. Visual captions: augmenting verbal communication with on-the-fly visuals. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–20.
- Livesley and Bromley (1973) W John Livesley and Dennis Basil Bromley. 1973. Person perception in childhood and adolescence. (1973).
- Logeswaran and Bhattacharya (2009) Nidhya Logeswaran and Joydeep Bhattacharya. 2009. Crossmodal transfer of emotion by music. Neuroscience letters 455, 2 (2009), 129–133.
- Mackey et al. (2020) Angella Mackey, Ron Wakkary, Stephan Wensveen, Annika Hupfeld, and Oscar Tomico. 2020. Alternative Presents for Dynamic Fabric. In Proceedings of the 2020 ACM Designing Interactive Systems Conference (DIS ’20). ACM. https://doi.org/10.1145/3357236.3395447
- Martey and Consalvo (2011) Rosa Mikeal Martey and Mia Consalvo. 2011. Performing the looking-glass self: Avatar appearance and group identity in Second Life. Popular Communication 9, 3 (2011), 165–180.
- Masullo et al. (2021) Massimiliano Masullo, Luigi Maffei, Tina Iachini, Mariachiara Rapuano, Federico Cioffi, Gennaro Ruggiero, and Francesco Ruotolo. 2021. A questionnaire investigating the emotional salience of sounds. Applied Acoustics 182 (2021), 108281.
- May and Hamilton (1980) James L May and Phyllis Ann Hamilton. 1980. Effects of musically evoked affect on women’s interpersonal attraction toward and perceptual judgments of physical attractiveness of men. Motivation and Emotion 4 (1980), 217–228.
- McGookin and Brewster (2012) David McGookin and Stephen Brewster. 2012. PULSE: the design and evaluation of an auditory display to provide a social vibe. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems. 1263–1272.
- McGookin et al. (2009) David McGookin, Stephen Brewster, and Pablo Priego. 2009. Audio bubbles: Employing non-speech audio to support tourist wayfinding. In Haptic and Audio Interaction Design: 4th International Conference, HAID 2009 Dresden, Germany, September 10-11, 2009 Proceedings 4. Springer, 41–50.
- Miller (1997) Kimberly A Miller. 1997. Dress: Private and secret self-expression. Clothing and Textiles Research Journal 15, 4 (1997), 223–234.
- Mueller and Karau (2002) Florian Mueller and Matthew Karau. 2002. Transparent hearing. In CHI’02 extended abstracts on Human Factors in Computing Systems. 730–731.
- Mynatt et al. (1998) Elizabeth D Mynatt, Maribeth Back, Roy Want, Michael Baer, and Jason B Ellis. 1998. Designing audio aura. In Proceedings of the SIGCHI conference on Human factors in computing systems. 566–573.
- Nacke et al. (2010) Lennart E Nacke, Mark N Grimshaw, and Craig A Lindley. 2010. More than a feeling: Measurement of sonic user experience and psychophysiology in a first-person shooter game. Interacting with computers 22, 5 (2010), 336–343.
- Oxford English Dictionary (2023) Oxford English Dictionary. 2023. persona (n.), sense 2.a. https://doi.org/10.1093/OED/1145536616
- Pijanowski et al. (2011) Bryan C Pijanowski, Almo Farina, Stuart H Gage, Sarah L Dumyahn, and Bernie L Krause. 2011. What is soundscape ecology? An introduction and overview of an emerging new science. Landscape ecology 26 (2011), 1213–1232.
- Rantala et al. (2018) Inka Rantala, Ashley Colley, and Jonna Häkkilä. 2018. Smart Jewelry: Augmenting Traditional Wearable Self-Expression Displays. In Proceedings of the 7th ACM International Symposium on Pervasive Displays (PerDis ’18). ACM. https://doi.org/10.1145/3205873.3205891
- Rios et al. (2018) Juan Sebastian Rios, Daniel John Ketterer, and Donghee Yvette Wohn. 2018. How Users Choose a Face Lens on Snapchat. In Companion of the 2018 ACM Conference on Computer Supported Cooperative Work and Social Computing (CSCW ’18). ACM. https://doi.org/10.1145/3272973.3274087
- Roinesalo et al. (2016) Paula Roinesalo, Juho Rantakari, Lasse Virtanen, and Jonna Häkkilä. 2016. Clothes integrated visual markers as self-expression tool. In Proceedings of the 18th International Conference on Human-Computer Interaction with Mobile Devices and Services Adjunct (MobileHCI ’16). ACM. https://doi.org/10.1145/2957265.2961832
- Rozier (2000) Joseph Michael Rozier. 2000. Here&There: an augmented reality system of linked audio. Ph. D. Dissertation. Massachusetts Institute of Technology.
- Schultze and Leahy (2009) Ulrike Schultze and Matthew Michael Leahy. 2009. The avatar-self relationship: Enacting presence in second life. (2009).
- Segalin et al. (2017) Cristina Segalin, Fabio Celli, Luca Polonio, Michal Kosinski, David Stillwell, Nicu Sebe, Marco Cristani, and Bruno Lepri. 2017. What your Facebook profile picture reveals about your personality. In Proceedings of the 25th ACM international conference on Multimedia. 460–468.
- Siddiqui (2021) Adeeba Siddiqui. 2021. Social media and its role in amplifying a certain idea of beauty. Infotheca—Journal for Digital Humanities 21, 1 (2021), 73–85.
- Siibak (2009) Andra Siibak. 2009. Constructing the self through the photo selection-visual impression management on social networking websites. Cyberpsychology: Journal of psychosocial research on cyberspace 3, 1 (2009), 1.
- Sikora et al. (2018) Marjan Sikora, Mladen Russo, Jurica Đerek, and Ante Jurčević. 2018. Soundscape of an archaeological site recreated with audio augmented reality. ACM Transactions on Multimedia Computing, Communications, and Applications (TOMM) 14, 3 (2018), 1–22.
- Spence (2013) Charles Spence. 2013. Just how important is spatial coincidence to multisensory integration? Evaluating the spatial rule. Annals of the New York Academy of Sciences 1296, 1 (2013), 31–49.
- Storek et al. (2015) Dominik Storek, Jan Stuchlik, and Frantisek Rund. 2015. Modifications of the surrounding auditory space by augmented reality audio: Introduction towarped acoustic reality. (2015).
- Swapan et al. (2019) Abu Yousuf Swapan, Joo Hwa Bay, and Dora Marinova. 2019. Importance of the residential front yard for social sustainability: Comparing sense of community levels in semi-private-public open spaces. Journal of Green Building 14, 2 (2019), 177–202.
- Tajadura and Västfjäll (2008) Ana Tajadura and Daniel Västfjäll. 2008. Auditory-induced emotion-A neglected channel for communication in HCI. Emotion in HCI (2008), 81.
- Tajadura-Jiménez et al. (2015a) Ana Tajadura-Jiménez, Maria Basia, Ophelia Deroy, Merle Fairhurst, Nicolai Marquardt, and Nadia Bianchi-Berthouze. 2015a. As light as your footsteps: altering walking sounds to change perceived body weight, emotional state and gait. In Proceedings of the 33rd annual ACM conference on human factors in computing systems. 2943–2952.
- Tajadura-Jiménez et al. (2011) Ana Tajadura-Jiménez, Galini Pantelidou, Pawel Rebacz, Daniel Västfjäll, and Manos Tsakiris. 2011. I-space: the effects of emotional valence and source of music on interpersonal distance. PloS one 6, 10 (2011), e26083.
- Tajadura-Jiménez et al. (2015b) Ana Tajadura-Jiménez, Manos Tsakiris, Torsten Marquardt, and Nadia Bianchi-Berthouze. 2015b. Action sounds update the mental representation of arm dimension: contributions of kinaesthesia and agency. Frontiers in psychology 6 (2015), 689.
- Tajadura-Jiménez et al. (2018) Ana Tajadura-Jiménez, Aleksander Väljamäe, Frédéric Bevilacqua, and Nadia Bianchi-Berthouze. 2018. Principles for Designing Body-Centered Auditory Feedback. The Wiley handbook of human computer interaction 1 (2018), 371–403.
- Tajfel (1974) Henri Tajfel. 1974. Social identity and intergroup behaviour. Social science information 13, 2 (1974), 65–93.
- Takayama and Nass (2010) Leila Takayama and Clifford Nass. 2010. Throwing voices: The psychological impact of the spatial height of projected voices. In Proceedings of the 2010 ACM conference on Computer supported cooperative work. 91–94.
- Todorov et al. (2013) Alexander Todorov, Ron Dotsch, Jenny M Porter, Nikolaas N Oosterhof, and Virginia B Falvello. 2013. Validation of data-driven computational models of social perception of faces. Emotion 13, 4 (2013), 724.
- Tonetto et al. (2014) Leandro Miletto Tonetto, Cristiano Porto Klanovicz, and Charles Spence. 2014. Modifying action sounds influences people’s emotional responses and bodily sensations. i-Perception 5, 3 (2014), 153–163.
- Turkle (2011) Sherry Turkle. 2011. Life on the Screen. Simon and Schuster.
- Turmo Vidal et al. (2024) Laia Turmo Vidal, José Manuel Vega-Cebrián, María Concepción Valdez Gastelum, Elena Márquez Segura, Judith Ley-Flores, Joaquin R Diaz Duran, and Ana Tajadura-Jiménez. 2024. Body Sensations as Design Material: An Approach to Design Sensory Technology for Altering Body Perception. In Proceedings of the 2024 ACM Designing Interactive Systems Conference. 2545–2561.
- Viinikainen et al. (2012) Mikko Viinikainen, Jari Kätsyri, and Mikko Sams. 2012. Representation of perceived sound valence in the human brain. Human brain mapping 33, 10 (2012), 2295–2305.
- Wang et al. (2017) Yanan Wang, Shijian Luo, Yujia Lu, Hebo Gong, Yexing Zhou, Shuai Liu, and Preben Hansen. 2017. AnimSkin: Fabricating Epidermis with Interactive, Functional and Aesthetic Color Animation. In Proceedings of the 2017 Conference on Designing Interactive Systems (DIS ’17). ACM. https://doi.org/10.1145/3064663.3064687
- Watanabe and Terada (2020) Hiroki Watanabe and Tsutomu Terada. 2020. Manipulatable auditory perception in wearable computing. In Proceedings of the Augmented Humans International Conference. 1–12.
- Weiser and Brown (1996) Mark Weiser and John Seely Brown. 1996. Designing calm technology. PowerGrid Journal 1, 1 (1996), 75–85.
- Won et al. (2015) Andrea Stevenson Won, Jeremy Bailenson, Jimmy Lee, and Jaron Lanier. 2015. Homuncular flexibility in virtual reality. Journal of Computer-Mediated Communication 20, 3 (2015), 241–259.
- Workman and Johnson (1991) Jane E Workman and Kim KP Johnson. 1991. The role of cosmetics in impression formation. Clothing and Textiles Research Journal 10, 1 (1991), 63–67.
- Yang et al. (2022) Jing Yang, Amit Barde, and Mark Billinghurst. 2022. Audio augmented reality: A systematic review of technologies, applications, and future research directions. journal of the audio engineering society 70, 10 (2022), 788–809.
- Yang et al. (2019) Jing Yang, Yves Frank, and Gábor Sörös. 2019. Hearing is believing: synthesizing spatial audio from everyday objects to users. In Proceedings of the 10th Augmented Human International Conference 2019. 1–9.
- Yang et al. (2020) Jing Yang, Prasanth Sasikumar, Huidong Bai, Amit Barde, Gábor Sörös, and Mark Billinghurst. 2020. The effects of spatial auditory and visual cues on mixed reality remote collaboration. Journal on Multimodal User Interfaces 14, 4 (2020), 337–352.
- Yang et al. (2018) Wanlu Yang, Kai Makita, Takashi Nakao, Noriaki Kanayama, Maro G Machizawa, Takafumi Sasaoka, Ayako Sugata, Ryota Kobayashi, Ryosuke Hiramoto, Shigeto Yamawaki, et al. 2018. Affective auditory stimulus database: An expanded version of the International Affective Digitized Sounds (IADS-E). Behavior Research Methods 50 (2018), 1415–1429.
- Yee and Bailenson (2007) Nick Yee and Jeremy Bailenson. 2007. The Proteus effect: The effect of transformed self-representation on behavior. Human communication research 33, 3 (2007), 271–290.
- Zaki (2013) Jamil Zaki. 2013. Cue integration: A common framework for social cognition and physical perception. Perspectives on Psychological Science 8, 3 (2013), 296–312.
- Zentner et al. (2008) Marcel Zentner, Didier Grandjean, and Klaus R Scherer. 2008. Emotions evoked by the sound of music: characterization, classification, and measurement. Emotion 8, 4 (2008), 494.
- Zhao et al. (2023) Yiran Zhao, Yujie Tao, Grace Le, Rui Maki, Alexander Adams, Pedro Lopes, and Tanzeem Choudhury. 2023. Affective Touch as Immediate and Passive Wearable Intervention. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies 6, 4 (2023), 1–23.

### v2

@@
 To instantiate the concept, we implemented a headphone-based prototype with multi-user tracking and audio streaming.
-Our formative study with designers revealed that audio personas were preferred in public and semi-public-private spaces for managing social impressions (e.g., personality) and signaling current states (e.g., emotions).
 Our preregistered in-lab study with 64 participants showed that audio personas influenced how participants formed impressions.
 Individuals with positive audio personas were rated as more socially attractive, more likable, and less threatening than those with negative audio personas.
+Our study with audio designers revealed that audio personas were preferred in public and semi-public-private spaces for managing social impressions (e.g., personality) and signaling current states (e.g., emotions).111Manuscript accepted to ACM Transactions on Computer-Human Interaction (TOCHI).
+This is the authors’ preprint version.
 1. Introduction
@@
 While people can easily modify their social images online, face-to-face interactions lack such flexibility.
-Physical adornments like makeup and clothing remain relatively static, limited in their ability convey changing social messages.
+Physical adornments like makeup and clothing remain relatively static, limited in their ability to convey changing social messages.
 As such, there is a growing trend toward integrating digital elements, traditionally limited to virtual interactions, into real-world settings.
@@
 In this work, we explore the opportunities to enhance face-to-face interactions through the auditory channel.
-We introduce the concept of audio personas111Here, a persona refers to the aspect of someone’s character that is presented to or perceived by others (Oxford English Dictionary, 2023; Goffman, 2023)., body-anchored audio cues to augment social perception in face-to-face interactions.
-Sound effects are commonly used in online games to provide immediate feedback on user actions and enhance engagement (Liljedahl, 2011; Lipscombc, 2013; Kenwright, 2020).
-However, it’s potential use face-to-face social contexts is largely unexplored.
-In fact, psychology research has shown that mere exposure to emotion-evoking sounds can influence how people perceive the social attractiveness and emotional states of others in photographs (May and Hamilton, 1980; Hanser et al., 2015; Logeswaran and Bhattacharya, 2009).
+We introduce the concept of audio personas, body-anchored audio cues to augment social perception in face-to-face interactions.
+Here, the term persona refers to the aspect of someone’s character that is presented to others (Jung, 2014; Goffman et al., 2002), which can vary depending on the social context (Jung, 2014; Goffman et al., 2002; Leary and Kowalski, 1990; Tajfel, 1974).
+Sound effects are commonly used in online games to provide immediate feedback for user actions like hitting a barrier, or collecting a reward, enhancing player engagement (Liljedahl, 2011; Lipscombc, 2013; Kenwright, 2020).
+However, using audio in face-to-face social contexts is largely unexplored.
+In fact, psychological studies showed that exposure to emotion-evoking sounds can affect how people perceive others’ social attractiveness and emotions (May and Hamilton, 1980; Hanser et al., 2015; Logeswaran and Bhattacharya, 2009).
+For instance, women rated photos of men as more attractive when listening to positive music compared to when they heard negative music (May and Hamilton, 1980).
 With the rise of spatial audio, the ability to spatially anchor sounds to individuals—similar to how visual filters personalize appearance—creates a novel means of personalized expression through sounds.
-Figure 1 visualizes a few examples of audio personas: one can choose to surround themselves with bee sounds to set personal boundaries, a ”woosh” sound to mimic passing by like a breeze for fun and playfulness, or keyboard-clicking sounds to signal that they are in a busy state.
+Figure 1 visualizes a few examples of audio personas.
+For instance, one can choose buzzing bee sounds to indicate personal boundaries, or a ”woosh” sound to mimic passing by like a breeze for fun and playfulness.
+One could also customize their footsteps to sound like bouncing balls or add a third footstep to express their love for gaming.
+Alternatively, keyboard-clicking sounds could be used to signal that the user is busy.
 To instantiate this concept, we first built a headphone-based prototype that utilizes multi-person tracking and multi-channel audio streaming.
 The user anchors sounds to their body and projects them to others in audio augmented reality.
-We conducted two user studies to explore the concept of audio personas from two perspectives: (1) when users design and project audio personas, i.e., senders, and (2) when users listen to the audio personas of others, i.e., receivers.
-Study #1 involved 8 designers to gain an initial understanding of the social intentions and contexts for using audio personas.
-We identified intended use cases for audio personas in public and semi-public-private spaces as a way to manage social impressions (e.g., personality) and signal current states (e.g., emotions).
-We also elicited the creations of 48 audio persona designs and found that audio personas being body-anchored is a key design feature.
-Additionally, we identified diverse design patterns in localization, triggers, dynamics and content of audio personas.
-Study #2 built upon the identified use cases, examining how audio personas influence social impression and perceived emotional states from the receiver’s perspective.
-In this preregistered in-lab study (n = 64), we focused on key design factors of audio personas (i.e., body-anchored, audio valence).
-Participants interacted with two research assistants, while hearing the research assistants’ audio personas or hearing sounds anchored to objects in the space.
-We found that participants were more likely to incorporate audio personas, body-anchored audio cues, into their impressions of others compared to the object-anchored audio.
+We conducted two user studies to explore the concept of audio personas.
+Study #1 (n = 64) examined whether audio personas are effective in influencing social perception, using a controlled, in-lab setup.
+Audio personas are designed to be body-anchored; therefore, as a between-subjects factor, we varied the spatial anchoring of sounds to assess the impact of this design choice.
+Participants interacted with two research assistants while hearing sounds either anchored to the assistants (i.e., audio personas) or to objects in the environment.
+As a within-subjects factor, we varied the valence of sounds to investigate how the emotional tone of audio personas influenced social perception.
+Overall, we found that participants were more likely to incorporate audio personas, body-anchored audio cues, into their impressions of others compared to the object-anchored audio.
 Moreover, participants perceived the person with a positive audio persona as more socially attractive, likable and less threatening than the person with a negative audio persona.
-These findings align with the theoretical grounding that social perception is a multisensory process (Livesley and Bromley, 1973; Workman and Johnson, 1991).
-The spatial alignment of audio with a person, as introduced by audio personas, enhances the likelihood of the sound being perceived as relevant cues of that individual in social interactions.
+As Study #1 demonstrated that audio personas can influence social perception, Study #2 aimed to gain an understanding of use cases and design possibilities for audio personas.
+Through a design study with 8 audio designers, we identified intended use cases of audio personas in public and semi-public-private spaces as a way to manage social impressions (e.g., personality) and signal current states (e.g., emotions).
+We also elicited the creations of 48 audio personas from the designer participants and identified diverse design patterns in localization, triggers, dynamics and content of audio personas.
 Our key contributions are:
@@
 The concept opens up possibilities for engaging dynamic, multisensory cues to shape social perception.
-Through a formative study with designers, we explored the social intentions and contexts for audio persona usage.
-This study also uncovered design possibilities for audio personas and identified key design patterns.
 In a preregistered in-lab study (n = 64), we showed that audio personas can influence how one forms impressions of others.
 These impressions were shaped in alignment with the valence of the audio persona.
+Through a design study with audio designers (n = 8), we explored the social intentions and contexts for audio persona usage.
+This study also uncovered design possibilities for audio personas and identified key design patterns.
+We discussed insights from the user studies and laid out design considerations and directions for future explorations of audio personas.
 2. Related Work
@@
 introduced the concept of ”sonic self-avatar,” emphasizing that self-representations in mixed reality environments should incorporate body-centered sounds (Tajadura-Jiménez et al., 2018; Tajadura and Västfjäll, 2008).
-When these body-centered sounds are altered, they can influence users’ mental representations of their bodies (Tajadura-Jiménez et al., 2015b; D’Adamo et al., 2024; Tonetto et al., 2014), perception of objects (Furfaro et al., 2015), and even gaits and motion (Tajadura-Jiménez et al., 2015a).
-For example, amplifying the high frequencies of footstep sound in walking can create the self-perception of having a thinner body (Tajadura-Jiménez et al., 2015a).
+When body actions are sonified, the sonification can influence users’ mental representations of their bodies (Tajadura-Jiménez et al., 2015b; D’Adamo et al., 2024; Tonetto et al., 2014; Frid et al., 2016), perception of objects (Furfaro et al., 2015), and even gaits and motion (Tajadura-Jiménez et al., 2015a).
+For example, amplifying the high frequencies of footstep sound can create the self-perception of having a thinner body (Tajadura-Jiménez et al., 2015a) and sonifying body movement with descending pitch increases movement acceleration (Singh et al., 2024).
 These audio-based sensory feedback devices were further explored in the context of physical activity and rehabilitation (Ley-Flores et al., 2024; Turmo Vidal et al., 2024; Gomez-Andres et al., 2020).
@@
 In the current setup, the maximum number of users the system can support is limited by the number of output devices the host computer can connect to.
-4. Study #1: User Perception and Designs of Audio Personas
-What do people want to express with audio personas, and how do they use audio persona to convey information?
-To explore user perception and the design possibilities of audio personas, we conducted a formative study.
-We chose to study participants with prior professional experience in audio.
-This sample selection allows us to understand the potential contexts and intentions behind audio persona usage from a user perspective, as well as elicit designs with expert insights.
-4.1. Methods
-4.1.1. Participants
-We recruited 8 participants with prior experience in producing/working with audio content from diverse backgrounds: audio engineering (), VR/AR design (), music () and immersive arts ().
-Each participant had at least 2 years of experience in producing audio content and received training from university-level programs.
-The participants were recruited through word-of-mouth.
-4.1.2. Procedure
-Participants were first introduced to the concept of audio personas.
-They put on the prototype system, as described in Sec.
-3, and interacted with an experimenter to experience three examples of audio personas: (1) a bee sound when approaching the experimenter, (2) a ”woosh” sound when the experimenter walked past, and (3) exaggerated stomping sounds from the experimenter.
-These demos intended to provide inspiration, and participants were encouraged to think beyond these audio persona designs.
-We then interviewed participants about the social contexts where they believe audio personas could be useful.
-We began with open-ended questions and then discussed audio persona usage in different types of social contexts—public (e.g., festivals, parks, streets, public transport), semi-public-private (e.g., workspace, educational institutions, community events), and private (e.g., home, private gathering) (Karaçor, 2016; Swapan et al., 2019).
-We also explored social messages suitable for audio personas to express.
-Again, we began with open-ended questions and discussed preferences for expressing specific types of messages using audio personas, including emotions, ideas and opinions, needs and desires, traits and personality, identity and belonging, and personal experiences (Green, 2007).
-Next, we elicited the creation of audio personas through two design sessions.
-Participants were asked to imagine themselves as designers of an audio persona library and create six different audio persona designs.
-In both sessions, participants used worksheets provided in the supplemental materials.
-In the first session, participants described what each audio persona should ”sound” like, its social contexts, and intended message.
-In the second design session, they specified sound characteristics, including spatial locations, triggers, and dynamics.
-At the end of each session, participants reviewed and discussed their designs with the experimenters.
-The study took approximately 75 - 90 minutes, and participants were compensated $35 for their time. The study was approved by the Institutional Review Board (IRB # 71992).
-4.1.3. Analysis
-We video-recorded the sessions, transcribed the data, and organized it based on the guiding interview questions, as described in the previous section and shown in Figure 3.
-Two experimenters participated in coding the transcriptions and designs, and we identified common themes through inductive thematic analysis (Braun and Clarke, 2006).
-The themes identified through thematic analysis are reflected in the headings of the following subsections.
-4.2. Finding #1: Social Intentions of Using Audio Personas
-Figure 4 shows participants’ preferences for using audio personas to express different types of social messages.
-All participants agreed that audio personas are well-suited for expressing ”personality & traits” and ”emotions”, followed by ”identity & belonging.” In contrast, only 2 out of 8 found audio personas suitable for expressing ”needs & desires” or ”personal experience,” with ”ideas & opinions” as the least favored.
-Together with qualitative findings, we identified two major social intentions for audio persona usage: managing social impressions (e.g., ”personality & traits,” ”identity & belonging”) and signaling current states (e.g., ”emotions”).
-We also identified the limitations of audio personas in conveying more complex information (e.g., ”needs & desires,” ”ideas & opinions,” ”personal experiences”).
-4.2.1. Impression management
-When discussing their intentions of using audio personas, P2, P3, P4 and P5 drew parallels with the selection of clothing.
-Similar to clothing, audio personas offer ”a way to express themselves” (P5), enabling others to ”get that first impression based on something that the other person set” (P2).
-Even more, P3 elaborated that audio persona are able to convey messages difficult to express through clothing:
-”Somebody’s audio persona would tell you another dimension of how they think …some of them might be kind of cheeky or humorous and you can’t necessarily or harder to convey with clothing.
-Unless you buy one of those T-shirts that have a joke on it; and those get really old really quickly.”
-As a dynamic medium, audio personas could serve as ”one of their intros when they’re meeting people” (P1), like a ”business cards” (P6), showing ”what they like, maybe what they don’t like” (P4), and fostering ”affinities” (P6).
-Moreover, audio personas were viewed as a means for creative expression.
-P8 described:
-”This is a nice way to kind of bring your personality to something more fictional, less rooted in reality …you can really customize what other’s perception of you is …I feel like your physical presence feels different because of the sound.”
-In the designs created by participants, 21 out of 48 audio personas intended to manage impressions, focusing on influencing how others perceive users.
-A wide range of audio content with varying valence was used and Figure 5 shows a selection of them.
-For instance, P7 designed an audio persona that features auditory accessories with a jingling bell sound to reflect personal jewelry taste and personality, as shown in (Figure 5a).
-P5 chose to augment their footstep sounds with a trombone sound, showing the love of SpongeBob (Figure 5b).
-P4 designed an audio persona that emits a dog barking sound whenever they carry their dog in a pouch, signaling that they are a pet owner and reminding others of the dog’s presence (Figure 5e).
-4.2.2. Signaling current states
-Another social intention for using audio personas is to signal the users’ current states and inform others on how they prefer to be interacted with.
-Audio personas offer an ”implicit” channel (Weiser and Brown, 1996) to convey current states, without saying out loud and ”imposing too much on [others]” (P4).
-P1 described it as ”another social cue in terms of emotions” and P5 saw it similar to ”setting your status” commonly found in chat apps.
-P7 described examples of how they might use audio personas to signal current states. Different valences of sounds were leveraged to signal different emotional states.
-”If I’m not feeling super chatty and I don’t want to engage in conversation with them, the [audio persona] can be something softer or more melodic, less exciting, more just like a kind of monotone.
-Versus like, if I’m feeling really high energy, it can be a more high energy sound”, said P7.
-Such indications of one’s current state could help others to ”empathize better with where one’s at” (P3) and ”help other people to know how to interact with you” (P8). For instance, P2 described:
-”If someone’s like, kind of down or depressed, like, we hear the rainfall?… maybe we could like, approach them and be like, ‘Oh, are you okay?’”
-21 out of 48 designs were created to signal current states, reflecting users’ emotions, needs, or intended interactions, and Figure 6 shows a selection of them.
-For instance, P3 used the Mission Impossible theme song to signal urgency in settings like a grocery store, prompting others to step aside or assist (Figure 6d).
-In contrast, P4 designed peaceful forest sounds to indicate serenity and a desire not to be disturbed during activities like meditation (Figure 6b).
-P6 created a bombastic drum and metal guitar persona to express their dislike for chores such as dishwashing (Figure 6c).
-4.2.3. Limitations in delivering complex information
-When discussing messages that audio personas might not be suitable to convey, participants referred to those involving complex information or narratives, such as opinions and personal experiences.
-P2 attributed the limitations to how audio is prone to ”perceptual bias” and P5 stated these information need ”more clarity”.
-For instance, when considering ideas and opinions, P8 described them as ”very nuanced”.
-”While audio is nuanced, it is always subjective.
-And so I don’t think it would clearly convey whatever idea or opinion would want it to pitch or rather, it has the potential to be misinterpreted”, said P8.
-Similarly, P4 expressed that personal experience is ”too complex a subject almost to capture in an audio clip”:
-”That’s like you have to imply something that is reminiscent of your childhood or reminiscent of a recent life experience and hope that people capture that implied message; but I don’t think it’ll be as effective as compared to some of the other options.”
-P1 suggested a good audio persona should be ”something that is not too long, like something that’s iconic, and then you could identify quickly.”
-4.3. Finding #2: Social Context for Using Audio Personas
-Figure 7 shows the participants’ rankings of their preferred social context for using audio personas.
-Public spaces (e.g., festivals, parks, streets, public transport) and semi-public-private spaces (e.g., workspace, educational institutions, community events) were found to be most preferred.
-Specifically, 6 out of 8 participants rated public spaces as where audio personas will be most useful for, followed by semi-public-private spaces.
-7 out of 8 participants found the private space to be the least useful for using audio personas.
-In terms of audio persona designs, 36 designs were created for semi-public-private environments, 14 for public spaces, and 5 for private settings.
-The total exceeds 48 because some designs are intended for multiple contexts.
-4.3.1. Intended for public and semi-public-private spaces
-In a public context, P6 described their experience of communications as:”you’re not actually able to verbalize your thoughts to as many people as you want as comfortably”.
-Thus, audio personas can provide a nice way to ”get first impressions of strangers” (P2), serve as ”a novel way to meet and greet” (P3) and ”communicate some sort of affinity” (P6).
-Thus, users can potentially ”make new connections” as audio personas become something that people can ”bond over” (P4).
-P3 echoed, ”In public contexts, audio personas could help you network with new people that are similar to you”.
-In semi-public-private spaces, P4 described the social interactions as: ”there are still gonna be a lot of strangers, and potentially coworkers.” Similarly, P2 described those spaces as more ”familiar” and they would ”become more open to different [audio] personas”.
-Similar to its use in public spaces, audio personas can serve as ”an interesting icebreaker …to get to know somebody else.” (P8).
-The usage of audio personas in semi-public-private space also involve ”understanding the vibes of somebody a little bit better” (P3).
-They elaborated with an example: ”When somebody is really zoned into something, they don’t want to be bothered at all …I can imagine, audio personas …being used for things like personal space violation.”
-4.3.2. Less useful in private space
-When considering a private social context, participants found it to be the least suitable for using audio personas.
-P1 described private gatherings as being: ”already pretty small.
-[and people are] in the same space and talking with each other.” P3 added that in those contexts, people usually ”know each other well enough”.
-Thus, participants deemed people should be able to ”start up a conversation” (P2), ”just talk to them honestly” (P4) and ”be fully present engaging with the person” (P8).
-In these contexts, audio personas were described as ”kind of unnecessary complication and distraction” (P3), making them feel ”less personal” (P5) and ”more disconnected from the people around them for conversation” (P7).
-On the other hand, P8 and P6 also described scenarios where they might want to use audio personas in a private context.
-P8 said: ”If I want to use it in a private space, it’s more of a novelty.” P6 envisioned an asynchronous usage where users re-experience other’s audio personas after private gatherings.
-4.4. Finding #3: Design Patterns of Audio Personas
-Lastly, we analyzed design patterns in the sound characteristics of audio persona designs, specifically examining their localization, activation triggers, and dynamic properties.
-4.4.1. Localizations
-We traced the sketches of audio persona locations marked by participants on the design worksheets and overlaid them to identify design patterns, as shown in Figure 8.
-One key observation across all 48 designs was that designers consistently anchored the audio persona to the user’s body, aligning the sound’s spatial location with the person’s.
-There was variability in the specific body parts where the audio personas were attached.
-25 out of 48 designs were anchored to the upper body of the users, such as to the head (n = 11), hands (n = 4), and back (n = 3).
-15 designs were placed around the lower body of the users, predominantly on the feet (n = 11).
-In 12 of 48 designs, participants circled the whole body instead of specifying a center origin, indicating they intended the audio personas to be heard around the body without a precise origin.
-8 designs placed audio personas on multiple body parts.
-While most designs were intended for individual use, P6 and P8 suggested that multiple people could share a group audio persona when engaging in similar activities, with each person contributing a part.
-In these cases, the audio persona would be placed at the group’s location, rather than from a single individual.
-4.4.2. Triggers
-Regarding what triggers the start of audio personas (i.e., when audio personas become audible to others), proximity is the most commonly used, as shown in Figure 9.
-Out of 48 designs, 30 were set to trigger when the distance between two users reached a predefined threshold.
-While most designs were triggered when others approached, in one of P6’s designs, they chose to trigger the audio persona only when someone walked past the user, calling it ”exit music.” 7 designs used specific gestures to trigger actions, such as the ”OK” hand gesture (P1) or air drumming motions (P7).
-2 designs relied on gaze, like eye contact, as the trigger.
-8 designs were context-dependent, activating in specific situations or when users engaged in certain activities, such as in a lecture hall (P6) or when a dog was nearby (P4).
-Additionally, participants explored using voice, body orientation (e.g., the direction the user is facing), or manual activation by the user.
-8 designs explored a combination of multiple triggers.
-4.4.3. Dynamics
-Most audio personas designs from Study # 1 are dynamic, with evolving sound characteristics.
-11 out of 48 designs involve a complete change in audio content based on factors like proximity, gestures, or user preferences.
-In one of P1’s designs, the audio persona shares the user’s favorite music, which adapts to the user’s changing taste.
-Other changes in dynamics involve the volume (n = 9), length (n = 7), pitch (n = 3), and speed (n = 2) of audio.
-For instance, in P2’s design, rain sounds vary in intensity to reflect the mood, becoming lighter as the person feels better.
-In 10 out of the 48 designs, participants chose to keep the audio persona unchanged.
-When multiple people are present, the blending of their audio personas become possible.
-For example, in P8’s design, a group shares an audio persona (e.g., forest or river sounds) during a group meditation to create a collective sense of calm.
-Each user retains their own sound and, as more people join, additional sounds, like bird sounds, are introduced.
-4.5. Summary of Findings in Study #1
-In Study #1, we explored social context and intended social messages for using audio personas by working with designer participants.
-We identified potential use cases in managing social impressions (e.g., personality & traits, identity & belonging) and signaling current states (e.g., emotions) but are limited in communicating complex information (e.g., ideas & opinions).
-Participants preferred using audio personas in public (e.g., street, transport) and semi-private settings (e.g., workspace, educational institutions) to form bonds with people they do not know well, but found it less useful in private spaces.
-Analysis of 48 participant-generated audio personas revealed that anchoring sounds to the body is a key design feature.
-Additionally, a wide range of audio content with varied valence was used, and there are diverse design possibilities for localization, triggers, and dynamics.
-5. Study #2: How Audio Persona Impacts Social Perception
-Study #1 uncovered the intended use of audio personas to manage social impressions and signal current states.
-To explore whether audio personas are effective in influencing these dimensions of social perception, we need to examine from the receiver’s perspective.
-Does listening to others’ audio personas change how one perceive others?
-In Study #2, 64 participants were invited to the lab for an icebreaker with two research assistants, with or without hearing the research assistants’ audio personas.
-As Study #1 identified body-anchored as a key design feature for audio persona and found that audio valence conveys distinct messages, Study #2 was designed around these two factors to examine the effect of audio personas.
-We hypothesized that hearing sounds anchored to a person (i.e., audio persona) would facilitate impression formation with the considerations of sounds.
-Furthermore, we hypothesized that the valence of the sounds would shape the direction of impressions and the perceived emotional states of others.
+4. Study #1: Can Audio Personas Affect Social Perception?
+To explore whether audio personas are effective in influencing social perception, we conducted a preregistered study to examine key design components of audio personas—specifically audio anchoring and audio valence—in a controlled setting.
+64 participants were invited to the lab to take part in an icebreaker activity with two research assistants.
 4.1. Condition Design
@@
 4.1.1. Between-group: audio anchoring.
-Audio personas being body-anchored was a key design feature observed in Study #1.
-Therefore, to understand the downstream impact of audio personas, it’s essential to understand how this spatial anchoring plays a role.
-Specifically, participants were divided into two groups.
-Half of the participants were assigned to the audio persona group (Figure 10a, 10b), in which participants heard sounds anchored to each research assistants.
+Audio personas are proposed to be body-anchored. Thus, to understand the effect of audio personas on social perception, it’s essential to understand how this spatial anchoring plays a role.
+Participants were divided into two groups.
+Half of the participants were assigned to the audio persona group (Figure 3a, 3b), in which participants heard sounds anchored to each research assistant.
 As the research assistant walked past participants (e.g., moving from left to right or vice versa), their audio persona shifted in locations corresponding to their movements.
@@
 We chose to include sounds in the control group as previous research showed that the mere presence of audio can already influence social emotional perception and behavior (Tajadura-Jiménez et al., 2011; Logeswaran and Bhattacharya, 2009; Carroll and Young, 2005).
-This control design thus better examine the exact effect of ”body-anchored” audio.
+This control design thus better examines the exact effect of ”body-anchored” audio.
 Anchoring sound to objects is also commonly used in audio augmented reality (McGookin et al., 2009; Yang et al., 2019; Sikora et al., 2018) and the auditory experience in this group is similar to hearing sounds played from a physical speaker.
@@
 4.1.2. Within-group: audio valence.
-In Study #1, participants leveraged sounds with varying valences in audio persona designs to tailor different contexts, aligning with prior research that showed audio valence can evoke distinct emotions in listeners (Zentner et al., 2008; Juslin and Västfjäll, 2008; Tajadura and Västfjäll, 2008).
-Therefore, Study #2 also examined how the emotional component of the sounds used in audio personas influences social perception.
+Prior research has shown that audio valence can evoke distinct emotions in listeners (Zentner et al., 2008; Juslin and Västfjäll, 2008; Tajadura and Västfjäll, 2008).
+Therefore, Study #1 also examined how the emotional component of the sounds used in audio personas influences social perception.
 Specifically, participants heard different types of sounds when interacting with each of the two research assistants.
@@
 Nested factor: audio content.
-As Study #1 highlighted the wide range of potential audio content that could be used in audio persona designs, Study #2 considered the generalizability of the results by incorporating diverse sounds.
+To ensure generalizability of the results, we incorporated diverse sounds.
 Specifically, we included two different sounds for each valence condition (positive and negative), resulting in a collection of two positive and two negative sounds.
@@
 4.1.3. Counterbalancing
-We used a Latin Square design to counterbalance the audio content participants heard when each research assistant was present.
-This resulted in 4 audio options for one research assistant and 2 options for the other one (covering two valence levels).
-Furthermore, we counterbalanced the side of the room each research assistant entered, which determined the participant’s standing location as described in Sec.
-5.1.2.
-Therefore, the sample size needed to be a multiple of (audio option for research assistant #1) (audio option for research assistant #2) (side of room entering) (condition groups) = to account for all combinations across both groups.
+We counterbalanced the audio valence (positive vs.
+negative) and the audio content (two options per valence) that participants heard during interactions with each research assistant.
+Additionally, we counterbalanced the side of the room from which each research assistant entered, which determined the participant’s standing position (as described in Section 5.1.2).
+Therefore, the sample size needed to be a multiple of (audio valence) × (audio content per valence) × (entry sides) × (research assistants) × (audio anchoring) = 32, to account for all combinations across both groups.
 Additionally, we randomized the order in which each research assistant entered and the side of the room where each speaker was located.
@@
 In each session, research assistants entered the room individually, walking back and forth in front of the participants.
-The changes to each research assistant were always on their name tags to remain consistency across participants: one switched from a yellow name tag to blue, and the other from a blue name tag to yellow.
-5.2.3. Research assistants training
+The changes to each research assistant were always on their name tags to remain consistent across participants: one switched from a yellow name tag to blue, and the other from a blue name tag to yellow.
+4.2.3. Familiarization phase
+Prior to the icebreaker, all participants completed a familiarization phase.
+This was intended to get all participants comfortable wearing the system, and also to prevent the control group from mistakenly attributing object-anchored sounds to the person.
+During this phase, participants were asked to walk around the space and identify whether any sounds were present and where they originated.
+The sounds were configured exactly as they would be during the icebreaker, and the control group heard sounds as they approached the corresponding objects.
+The research assistant was not present during this phase to ensure that participants encountered the research assistant for the first time during the main session.
+As a result, the audio persona group did not hear any sounds during this phase.
+All participants were able to identify whether sounds were present and, if so, determine their sources before proceeding with the icebreaker.
+4.2.4. Research assistants
 All participants interacted with the same two male, Asian research assistants, who wore identical clothing across study sessions to minimize the influence of attire on social perception.
@@
 Both research assistants were trained to maintain neutral body language and facial expressions in the study, similar to (Tajadura-Jiménez et al., 2011).
-The research assistants were blinded about the conditions or sounds assigned in all sessions to ensure consistent behavior.
-5.2.4. Familiarization phase
-Before the icebreaker, participants walked around the study space to familiarize themselves.
-In the control group, participants heard sounds as they approached the speakers on each side of the room, as they would later during the task.
-In the audio persona group, they did not hear anything since no research assistants were present.
-Participants identified the presence of sounds and their sources before proceeding with the task.
+They were instructed to walk at a steady pace, avoid expressive gestures and facial expressions, and maintain forward-facing eye gaze.
+An experimenter provided real-time feedback during repeated practice to ensure consistent and neutral behavior.
+In the study sessions, research assistants were blinded about the conditions assigned to the participants.
 4.3. Apparatus
@@
 Both body-anchored and object-anchored sounds were configured with a linear roll-off, with a minimum distance of 0.5 meters and a maximum distance of 2 meters.
-A 2-meter trigger distance was chosen based on audio persona designs from Study #1.
+A 2-meter trigger distance was chosen following E.T.
+Hall’s ”social distance” range (Hall, 1963).
 Specifically, the audio persona group began hearing sound from a distance of 2 meters, with the volume increasing as the research assistants moved closer to them.
@@
 Before starting the task, researchers asked participants to walk around the study space to become familiar with it.
-Additionally, researchers showed participants the designated standing locations for observing the research assistants, as described in Sec.
-5.1.
+Additionally, researchers showed participants the designated standing locations for observing the research assistants, as described previously.
 The task included two observation sessions.
@@
 Once participants had shown their answers to both research assistants, they were asked to place their answer sheets back on the table and return to sit on one of the benches.
-The experimenter then entered to reveal the correct answer to the change and guided the participants to another room to complete a post-task questionnaire.
-The study was approved by the Institutional Review Board (IRB # 71992). The full study took approximately 30 minutes, and all participants were compensated $15 for their time.
+During this phase, behavioral measures including participants’ physical distance from each research assistant and their seat selection were recorded.
+Participants could still hear the sounds anchored to the person or object if they were within the trigger distance.
+At the end, experimenters entered to reveal the correct answer to the change and guided the participants to another room to complete a post-task questionnaire.
+The study was approved by Stanford Institutional Review Board (IRB # 71992). The full study took approximately 30 minutes, and all participants were compensated $15 for their time.
 | Construct | Question |
@@
 4.5. Measures
-We measured social impression and perceived current states of others, two major uses of audio personas identified in Study #1.
-We collected both self-report and behavioral data.
-In this manuscript, we focus on discussing the self-report measures, and all behavioral measures details are provided in the supplemental materials.
-To understand how audio personas shape how one forms impressions of others, we started with impression drawings and impression writings.
+We adopted both self-report and behavioral measures to understand how audio personas might affect social perception.
+4.5.1. Self-report measures
+We started with impression drawings and impression writings.
 Drawing is a common technique in psychology to measure social perception (Kamphaus and Pleiss, 1991; Cherney et al., 2006), and we used drawing to understand if participants formed impressions of others with explicit consideration of their audio persona.
@@
 In support of the drawings, participants also wrote down their impressions of each research assistant using at least five sentences, referred to as impression writings.
-Two researchers independently coded the drawings and writings to analyze if participants referenced sounds through verbal or visual representations and resolved discrepancies through discussions.
 To capture the nuances of social impressions beyond the incorporation of sounds, we included constructs such as likability judgments, threat perception, and interpersonal attraction, inspired by prior work (Fauville et al., 2022).
@@
 Interpersonal attraction was measured by combining both social attraction (e.g., “I would get along with this person”) and task attraction (e.g., “I would like to interact with the person again.”), resulting in a 4-item construct (Fauville et al., 2022; Davis and Perkowitz, 1979).
-To examine whether audio personas influence how people interpret the current states of others, we focused on the perceived emotions of others.
-Emotions were among the top-rated preferred messages participants would like to express using audio personas, as shown in Study #1.
+Moreover, we examined whether audio personas influence how people interpret the emotional states of others, as prior work has shown that music can shape the perceived emotions of individuals depicted in photographs.
+(Logeswaran and Bhattacharya, 2009; Hanser et al., 2015).
 Specifically, we used the I-PANAS-SF scale (Karim et al., 2011), which includes 5 positive (active, attentive, alert, determined, inspired) and 5 negative terms (hostile, ashamed, upset, afraid, nervous), shown in Table 1.
-We assessed reliability using Cronbach’s alpha for constructs with more than two items and Spearman-Brown formula for two-item constructs.
+4.5.2. Behavioral measures
+Two behavioral measures were collected during the icebreaker activity: interpersonal distance and seat choice.
+Interpersonal distance was determined by calculating the Euclidean distance between the head positions (x and z axes) of the participant and each research assistant.
+Head positions and rotations were continuously tracked throughout the study session using head-mounted trackers.
+Specifically, we focused on the minimum interpersonal distance when participants were instructed to walk toward each research assistant to present their answer sheets.
+This task approximates the active approach paradigms used in prior work (Tajadura-Jiménez et al., 2011).
+In addition, we recorded which bench participants chose to sit on at the end of the icebreaker (i.e., seat choice), with each research assistant sitting on one.
+The selection of a seat closer to one research assistant over the other reflects the degree of social closeness participants felt (Hayduk, 1983).
+4.5.3. Overview of analysis
+To analyze impression drawing and writings, two researchers independently coded the drawings and writings to analyze if participants referenced sounds through verbal or visual representations and resolved discrepancies through discussions.
+As exploratory measures, we analyzed the count of sound-related terms and sentiments in impression writings.
+For self-report rating measures, we first assessed reliability using Cronbach’s alpha for constructs with more than two items and Spearman-Brown formula for two-item constructs.
 Initial validation showed that reliability for likability (r) and threat perception (r) were lower than 0.7.
@@
 For the two-item threat perception construct, individual item results were reported, due to the inability to improve reliability via factor analysis.
+We then used mixed ANOVA with audio anchoring as a between-group factor and audio valence as a within-group factor to analyze each of the self-report rating measures.
+For the behavioral measures, we also used mixed ANOVA to analyze the effect on minimum interpersonal distance.
+For the seat choice measure, since participants chose between one of the two benches to sit, we used a chi-square test to examine the relationship between audio valence and seat choice in each condition group.
 4.6. Hypotheses
@@
 None of the participants knew the research assistants before the study.
+4.8. Overview of Findings
+We organized the study findings into four sections.
+Finding #1, Inclusion of Audio in Impression Formation, examines whether participants considered audio as part of their impression of others, manifested in the impression drawings and writings.
+Finding #2, Impact on Social Impression, presents self-reported measures related to social impressions (social attraction, perceived threat, likability judgments), and sentiment analysis of the impression writings.
+Finding #3, Impact on Perceived Emotional States, includes participants’ ratings on the perceived emotional states of others.
+Finding #4, Impact on Social Behavior, reports participants’ seat choice selection and interpersonal distance.
 4.9. Finding #1: Inclusion of Audio in Impression Formation
@@
 This indicates that participants generally evaluated the emotional states of both of their partners similarly, regardless of the audio valence or anchoring.
+4.12. Finding #4: Impact on Social Behavior
+Seat choice.
+We first analyzed the seat picking choices at the end of the icebreaker task.
+In the audio persona group, 14 out of 32 participants chose to sit with the research assistant that had a positive audio persona while 18 chose to sit with the research assistant with a negative audio persona.
+A chi-square test was conducted to examine the relationship between audio persona valence and choice of seat selection.
+The test statistic was , indicating that there is no statistically significant relationship between the valence of the audio persona and the participants’ choice of seating.
+Similarly, in the control group, no significant relationship was found between the valence of sounds anchored to the objects objects and participants’ final seat selections ().
+Specifically, 20 out of the 32 participants sat closer to the research assistant paired with a positive sound and 12 participants sat closer to the research assistant paired with a negative sound, with the sounds anchored to objects in the space.
+Minimum interpersonal distance.
+Similarly, we analyzed the minimum interpersonal distance as participants approached each research assistant to show their answer sheets.
+This distance was calculated using the Euclidean distance between their head positions.
+Mixed ANOVA analysis revealed no significant main effects of either audio anchoring (F) or audio valence () on the minimum interpersonal distance.
+This suggests that participants generally kept a consistent distance from their partners, regardless of whether they were exposed to positive or negative sounds, and whether the sound was anchored to people or objects in the space.
+4.13. Summary of Findings in Study #1
+Through a controlled in-lab study, we examined the effect of audio personas on social perception.
+Participants exposed to audio personas (i.e., body-anchored audio cues) actively integrated these sounds into their impressions of others, as evidenced by their impression writings and drawings.
+In contrast, participants in the control group—who heard the same sounds anchored to objects—were much less likely to associate the sounds with the person.
+This contrast highlights body-anchoring as a key design feature that enables audio personas to influence social perception.
+Furthermore, individuals with a positive audio persona were rated as more socially attractive, likable, and less threatening than those with a negative audio persona.
+Lastly, we found no significant effects of audio personas on perceived emotional states of others or on behavioral measures.
+5. Study #2: User Scenarios and Designs of Audio Personas
+As Study #1 demonstrated the effectiveness of audio personas in influencing social perception, Study #2 explores potential use cases and design patterns for the concept.
+We conducted a design study with eight audio designers to uncover (1) social intentions behind using audio personas, (2) social contexts to use audio personas, and (3) design patterns of audio personas.
+5.1. Methods
+5.1.1. Participants
+We chose to study participants with prior professional experience in audio design.
+This sample selection allows us to understand the potential contexts and intentions behind audio persona usage from a user perspective, as well as elicit designs with expert insights.
+We recruited 8 participants with prior experience in producing/working with audio content from diverse backgrounds: audio engineering (), VR/AR design (), music () and immersive arts ().
+Each participant had at least 2 years of experience in producing audio content and received training from university-level programs.
+The participants were recruited through word-of-mouth.
+5.1.2. Procedure
+Participants were first introduced to the concept of audio personas.
+They put on the prototype system, as described in Sec.
+3, and interacted with an experimenter to experience three examples of audio personas: (1) a bee sound when approaching the experimenter, (2) a ”woosh” sound when the experimenter walked past, and (3) exaggerated stomping sounds from the experimenter.
+We chose these three demonstrations to show the technical possibilities of our prototype system, such as triggering audio personas based on proximity and anchoring sounds to body parts.
+These demos intended to provide inspiration, and participants were encouraged to think beyond these demo designs.
+We then interviewed participants about the social contexts where they believe audio personas could be useful.
+We began with open-ended questions and then discussed audio persona usage in different types of social contexts—public (e.g., festivals, parks, streets, public transport), semi-public-private (e.g., workspace, educational institutions, community events), and private (e.g., home, private gathering) (Karaçor, 2016; Swapan et al., 2019).
+We also explored social messages suitable for audio personas to express.
+Again, we began with open-ended questions and discussed preferences for expressing specific types of messages using audio personas, including emotions, ideas and opinions, needs and desires, traits and personality, identity and belonging, and personal experiences (Green, 2007).
+Next, we elicited the creation of audio personas through two design sessions.
+Participants were asked to imagine themselves as designers of an audio persona library and create six different audio persona designs.
+In both sessions, participants used worksheets provided in the supplemental materials.
+In the first session, participants described what each audio persona should ”sound” like, its social contexts, and intended message.
+In the second design session, they specified sound characteristics, including spatial locations, triggers, and dynamics.
+At the end of each session, participants reviewed and discussed their designs with the experimenters.
+The study took approximately 75 - 90 minutes, and participants were compensated $35 for their time. The study was approved by Stanford Institutional Review Board (IRB # 71992).
+5.1.3. Analysis
+We video-recorded the sessions, transcribed the data, and organized it based on the guiding questions described previously.
+Two researchers were involved in the coding process.
+One researcher conducted an initial inductive thematic analysis of the transcripts to identify common themes (Braun and Clarke, 2006).
+The themes derived from the thematic analysis are reflected in the headings of the following subsections.
+Based on these themes, the two researchers independently coded each design created by the audio designers.
+They then discussed any discrepancies to reach consensus.
+5.2. Finding #1: Social Intentions of Using Audio Personas
+Figure 10 shows participants’ preferences for using audio personas to express different types of social messages.
+All participants agreed that audio personas are well-suited for expressing ”personality & traits” and ”emotions”, followed by ”identity & belonging.” In contrast, only 2 out of 8 found audio personas suitable for expressing ”needs & desires” or ”personal experience,” with ”ideas & opinions” as the least favored.
+Together with qualitative findings, we identified two major social intentions for audio persona usage: managing social impressions (e.g., ”personality & traits,” ”identity & belonging”) and signaling current states (e.g., ”emotions”).
+We also identified the limitations of audio personas in conveying more complex information (e.g., ”needs & desires,” ”ideas & opinions,” ”personal experiences”).
+5.2.1. Impression management
+When discussing their intentions of using audio personas, P2, P3, P4 and P5 drew parallels with the selection of clothing.
+Similar to clothing, audio personas offer ”a way to express themselves” (P5), enabling others to ”get that first impression based on something that the other person set” (P2).
+Even more, P3 elaborated that audio persona are able to convey messages difficult to express through clothing:
+”Somebody’s audio persona would tell you another dimension of how they think …some of them might be kind of cheeky or humorous and you can’t necessarily or harder to convey with clothing.
+Unless you buy one of those T-shirts that have a joke on it; and those get really old really quickly.”
+As a dynamic medium, audio personas could serve as ”one of their intros when they’re meeting people” (P1), like a ”business cards” (P6), showing ”what they like, maybe what they don’t like” (P4), and fostering ”affinities” (P6).
+Moreover, audio personas were viewed as a means for creative expression.
+P8 described:
+”This is a nice way to kind of bring your personality to something more fictional, less rooted in reality …you can really customize what other’s perception of you is …I feel like your physical presence feels different because of the sound.”
+In the designs created by participants, 21 out of 48 audio personas intended to manage impressions, focusing on influencing how others perceive users.
+A wide range of audio content with varying valence was used and Figure 11 shows a selection of them.
+For instance, P7 designed an audio persona that features auditory accessories with a jingling bell sound to reflect personal jewelry taste and personality, as shown in (Figure 11a).
+P5 chose to augment their footstep sounds with a trombone sound, showing the love of SpongeBob (Figure 11b).
+P4 designed an audio persona that emits a dog barking sound whenever they carry their dog in a pouch, signaling at a distance that they are a pet owner and cuing others to notice the dog’s presence (Figure 11e).
+5.2.2. Signaling current states
+Another social intention for using audio personas is to signal the users’ current states and inform others on how they prefer to be interacted with.
+Audio personas offer an ”implicit” channel (Weiser and Brown, 1996) to convey current states, without saying out loud and ”imposing too much on [others]” (P4).
+P1 described it as ”another social cue in terms of emotions” and P5 saw it similar to ”setting your status” commonly found in chat apps.
+P7 described examples of how they might use audio personas to signal current states. Different valences of sounds were leveraged to signal different emotional states.
+”If I’m not feeling super chatty and I don’t want to engage in conversation with them, the [audio persona] can be something softer or more melodic, less exciting, more just like a kind of monotone.
+Versus like, if I’m feeling really high energy, it can be a more high energy sound”, said P7.
+Such indications of one’s current state could help others to ”empathize better with where one’s at” (P3) and ”help other people to know how to interact with you” (P8). For instance, P2 described:
+”If someone’s like, kind of down or depressed, like, we hear the rainfall?… maybe we could like, approach them and be like, ‘Oh, are you okay?’”
+21 out of 48 designs were created to signal current states, reflecting users’ emotions, needs, or intended interactions, and Figure 12 shows a selection of them.
+For instance, P3 used the Mission Impossible theme song to signal urgency in settings like a grocery store, prompting others to step aside or assist (Figure 12d).
+In contrast, P4 designed peaceful forest sounds to indicate serenity and a desire not to be disturbed during activities like meditation (Figure 12b).
+P6 created a bombastic drum and metal guitar persona to express their dislike for chores such as dishwashing (Figure 12c).
+5.2.3. Limitations in delivering complex information
+When discussing messages that audio personas might not be suitable to convey, participants referred to those involving complex information or narratives, such as opinions and personal experiences.
+P2 attributed the limitations to how audio is prone to ”perceptual bias” and P5 stated these information need ”more clarity”.
+For instance, when considering ideas and opinions, P8 described them as ”very nuanced”.
+”While audio is nuanced, it is always subjective.
+And so I don’t think it would clearly convey whatever idea or opinion would want it to pitch or rather, it has the potential to be misinterpreted”, said P8.
+Similarly, P4 expressed that personal experience is ”too complex a subject almost to capture in an audio clip”:
+”That’s like you have to imply something that is reminiscent of your childhood or reminiscent of a recent life experience and hope that people capture that implied message; but I don’t think it’ll be as effective as compared to some of the other options.”
+P1 suggested a good audio persona should be ”something that is not too long, like something that’s iconic, and then you could identify quickly.”
+5.2.4. Additional use cases for audio personas.
+Six designs did not fall into impression management or signaling current states.
+Two were situated in motor learning contexts.
+For instance, P05 designed smoothing and relaxing sounds that played when the teacher’s and student’s body motions were aligned.
+In this case, the audio served as immediate feedback audible to both the sender and receiver.
+Two designs leveraged audio personas to evoke or recall past events.
+For example, P08 used the sound of ocean waves while retelling a vacation experience, immersing the listeners in the story.
+We also observed audio persona use in accessibility contexts.
+By listening to changes in one’s own audio personas during locomotion (P08), blind and low vision individuals could gain spatial awareness of the space.
+Last but not least, audio personas were used as motivational social gestures, allowing users to send encouragement—such as the message “you got this”—to others.
+5.3. Finding #2: Social Context for Using Audio Personas
+Figure 13 shows the participants’ rankings of their preferred social context for using audio personas.
+Public spaces (e.g., festivals, parks, streets, public transport) and semi-public-private spaces (e.g., workspace, educational institutions, community events) were found to be most preferred.
+Specifically, 6 out of 8 participants rated public spaces as where audio personas will be most useful for, followed by semi-public-private spaces.
+7 out of 8 participants found the private space to be the least useful for using audio personas.
+In terms of audio persona designs, 36 designs were created for semi-public-private environments, 14 for public spaces, and 5 for private settings.
+The total exceeds 48 because some designs are intended for multiple contexts.
+5.3.1. Intended for public and semi-public-private spaces
+In a public context, P6 described their experience of communications as:”you’re not actually able to verbalize your thoughts to as many people as you want as comfortably”.
+Thus, audio personas can provide a nice way to ”get first impressions of strangers” (P2), serve as ”a novel way to meet and greet” (P3) and ”communicate some sort of affinity” (P6).
+Thus, users can potentially ”make new connections” as audio personas become something that people can ”bond over” (P4).
+P3 echoed, ”In public contexts, audio personas could help you network with new people that are similar to you”.
+In semi-public-private spaces, P4 described the social interactions as: ”there are still gonna be a lot of strangers, and potentially coworkers.” Similarly, P2 described those spaces as more ”familiar” and they would ”become more open to different [audio] personas”.
+Similar to its use in public spaces, audio personas can serve as ”an interesting icebreaker …to get to know somebody else.” (P8).
+The usage of audio personas in semi-public-private space also involve ”understanding the vibes of somebody a little bit better” (P3).
+They elaborated with an example: ”When somebody is really zoned into something, they don’t want to be bothered at all …I can imagine, audio personas …being used for things like personal space violation.”
+5.3.2. Less useful in private space
+When considering a private social context, participants found it to be the least suitable for using audio personas.
+P1 described private gatherings as being: ”already pretty small.
+[and people are] in the same space and talking with each other.” P3 added that in those contexts, people usually ”know each other well enough”.
+Thus, participants deemed people should be able to ”start up a conversation” (P2), ”just talk to them honestly” (P4) and ”be fully present engaging with the person” (P8).
+In these contexts, audio personas were described as ”kind of unnecessary complication and distraction” (P3), making them feel ”less personal” (P5) and ”more disconnected from the people around them for conversation” (P7).
+On the other hand, P8 and P6 also described scenarios where they might want to use audio personas in a private context.
+P8 said: ”If I want to use it in a private space, it’s more of a novelty.” P6 envisioned an asynchronous usage where users re-experience other’s audio personas after private gatherings.
+5.4. Finding #3: Design Patterns of Audio Personas
+Lastly, we analyzed design patterns in the sound characteristics of audio persona designs, specifically examining their localization, activation triggers, and dynamic properties.
+5.4.1. Localizations
+We traced the sketches of audio persona locations marked by participants on the design worksheets and overlaid them to identify design patterns, as shown in Figure 14.
+There was variability in the specific body parts where the audio personas were attached.
+25 out of 48 designs were anchored to the upper body of the users, such as to the head (n = 11), hands (n = 4), and back (n = 3).
+15 designs were placed around the lower body of the users, predominantly on the feet (n = 11).
+In 12 of 48 designs, participants circled the whole body instead of specifying a center origin, indicating they intended the audio personas to be heard around the body without a precise origin.
+8 designs placed audio personas on multiple body parts.
+While most designs were intended for individual use, P6 and P8 suggested that multiple people could share a group audio persona when engaging in similar activities, with each person contributing a part.
+In these cases, the audio persona would be placed at the group’s location, rather than from a single individual.
+5.4.2. Triggers
+Regarding what triggers the start of audio personas (i.e., when audio personas become audible to others), proximity is the most commonly used, as shown in Figure 15.
+Out of 48 designs, 30 were set to trigger when the distance between two users reached a predefined threshold.
+While most designs were triggered when others approached, in one of P6’s designs, they chose to trigger the audio persona only when someone walked past the user, calling it ”exit music.” 7 designs used specific gestures to trigger actions, such as the ”OK” hand gesture (P1) or air drumming motions (P7).
+2 designs relied on gaze, like eye contact, as the trigger.
+8 designs were context-dependent, activating in specific situations or when users engaged in certain activities, such as in a lecture hall (P6) or when a dog was nearby (P4).
+Additionally, participants explored using voice, body orientation (e.g., the direction the user is facing), or manual activation by the user.
+8 designs explored a combination of multiple triggers.
+5.4.3. Dynamics
+Most audio personas designs from Study # 1 are dynamic, with evolving sound characteristics.
+11 out of 48 designs involve a complete change in audio content based on factors like proximity, gestures, or user preferences.
+In one of P1’s designs, the audio persona shares the user’s favorite music, which adapts to the user’s changing taste.
+Other changes in dynamics involve the volume (n = 9), length (n = 7), pitch (n = 3), and speed (n = 2) of audio.
+For instance, in P2’s design, rain sounds vary in intensity to reflect the mood, becoming lighter as the person feels better.
+In 10 out of the 48 designs, participants chose to keep the audio persona unchanged.
+When multiple people are present, the blending of their audio personas become possible.
+For example, in P8’s design, a group shares an audio persona (e.g., forest or river sounds) during a group meditation to create a collective sense of calm.
+Each user retains their own sound and, as more people join, additional sounds, like bird sounds, are introduced.
+5.5. Summary of Findings in Study #2
+In Study #2, we explored social context and intended social messages for using audio personas by working with participants experienced in audio design.
+We identified potential use cases in managing social impressions (e.g., personality & traits, identity & belonging) and signaling current states (e.g., emotions) but are limited in communicating complex information (e.g., ideas & opinions).
+Participants preferred using audio personas in public (e.g., street, transport) and semi-private settings (e.g., workspace, educational institutions) to form bonds with people they do not know well, but found it less useful in private spaces.
+Analysis of 48 participant-generated audio personas revealed that a wide range of audio content with varied valence was used, and there are diverse design possibilities for localization, triggers, and dynamics.
 6. Discussion
@@
 6.1. Sound, Impression Management and Social Perception
-In Study #1, we studied with designer-participants and identified potential use cases of audio personas to manage social impressions (e.g., personality), and signal current states (e.g., emotions).
-Study #2 examined the effectiveness of audio persona in influencing these dimensions from the receiver’s standpoint.
-Here, we discussed key insights from the findings of both studies.
-Study #1 revealed that body anchoring is a key feature for audio personas, consistently present in all 48 designs created by designer-participants.
-Study #2 further showed that such spatial anchoring of sound shaped how social impressions were formed.
-Participants explicitly integrate audio personas into their impressions of others, revealed in the impression writings and drawings.
-This finding was contrasted with control group, where participants heard the same sounds anchored to objects and were much less likely to integrate sounds into their impressions of others.
+Study #1 employed an in-lab, controlled study to examine the effectiveness of audio personas, body-anchored sounds, in influencing social perception.
+Our findings showed that participants integrate audio personas into their impressions of others, revealed in the impression writings and drawings.
+In the control group, where participants heard the same sounds anchored to objects, it was much less likely to integrate sounds into their impressions of others.
 These findings align with a theoretical framework of impression formation (Livesley and Bromley, 1973; Workman and Johnson, 1991), which suggests that individuals form impressions by focusing on specific cues from the available information about a person and making inferences based on those cues.
@@
 Such a multisensory integration process could explain impression formation in the context of audio personas.
-Study #1 also showed that participants leveraged different valences of sound to convey various messages.
-Study #2 also investigated the emotional component of audio persona.
+Study #1 also investigated the emotional component of audio persona.
 We found participants rated the person with a positive audio persona as more socially attractive, likable, and less threatening than the person with a negative audio persona.
@@
 While previous research suggested that sounds prime the perceived emotions of others (Logeswaran and Bhattacharya, 2009; Ignacio et al., 2021; Carroll and Young, 2005), our study did not observe a similar effect.
-One possible explanation is that our study put participants into more complex social settings (beyond observing static faces as in prior work), where interpreting perceived emotions of others involve processing other cues (Lange et al., 2022).
+One possible explanation is that our study put participants into more complex social settings (beyond observing static faces as in prior work), where interpreting perceived emotions of others involves processing other cues (Lange et al., 2022).
 Since most non-verbal cues were controlled and kept neutral in our study, participants likely perceived the primary signals as neutral, overriding the priming effect of audio.
 This indicates that semantic congruence (Laurienti et al., 2004; Kim and Lee, 2022) between audio personas and other contextual information might be necessary for influencing the perceived emotional states of others.
+We did not observe main effects on certain behavioral measures, such as seat choice and interpersonal distance.
+These results are partially consistent with prior research on how sound influences interpersonal distance.
+While previous studies have found effects on audio in interpersonal distance, these were typically observed in passive approach tasks—where the experimenter approached the participant, and the participant indicated when to stop (Tajadura-Jiménez et al., 2011).
+In contrast, no effects were found in conditions where participants actively approached the experimenter (Tajadura-Jiménez et al., 2011).
+In our study, participants actively chose where to position themselves, which may explain the lack of significant behavioral effects and need further investigation.
+6.2. Design Considerations for Audio Personas
+Study #2 involved audio designers to elicit potential use cases and design patterns for audio personas. Based on these findings, we synthesized following design considerations.
+Designing for shared spaces.
+Considering social contexts to use audio personas, participants in our design study preferred using audio personas in public (e.g., street, transport) and semi-private settings (e.g., workspace, educational institutions) to form bonds with people they do not know well.
+As audio personas are envisioned for these shared spaces, a key design challenge is how to effectively manage the projection and reception of multiple audio personas.
+It’s easy to imagine situations where hearing multiple audio personas simultaneously could become distracting or overwhelming.
+One potential solution is to draw inspiration from online social media platforms, where users manage their networks by choosing who to subscribe or unsubscribe.
+Similarly, audio persona systems could give users control over who can hear their audio persona and whose personas they want to receive.
+Additionally, as suggested by several designs in from Study #2, blending or mixing multiple personas into a single ambient layer could help users sense the overall ”vibe” of the space without being overwhelmed by different sounds.
+Users could then selectively tune into a specific person’s audio persona if desired.
+Privacy, security and ethical considerations.
+Considering the social intentions behind using audio personas, participants in our design study expressed a preference for using them to manage social impressions (e.g., personality & traits, identity & belonging) and signaling current states (e.g., emotions).
+Building upon these use cases, security, privacy and ethical considerations need to be taken into account.
+Given that audio personas may contain personal information intended for others, privacy protections are essential.
+Users should have full control over with whom their audio personas are shared.
+It is also important that users are made aware of when their audio personas are being heard during interactions—for example, through notifications.
+Additionally, because audio personas are tied to a user’s spatial location, continuous broadcasting may unintentionally reveal behavioral patterns that users did not intend to share.
+To mitigate this risk, designers might consider suggesting spatial and temporal limits—such as setting the maximum distance within which an audio persona can be heard, or maximum duration it remains active.
+Designers should also be mindful of the potential harms associated with misuse of audio personas and develop appropriate intervention strategies.
+For instance, malicious actors could exploit the system by broadcasting loud, overwhelming sounds to capture attention or by transmitting inappropriate audio content.
+To mitigate such risks, one approach is to implement content moderation mechanisms, such as pre-filtering, to block disruptive or harmful audio before it reaches listeners.
+Additionally, in cases where audio personas are also audible to the senders themselves, potential effects on gait and movement should be considered, as highlighted by prior work on movement sonification (D’Adamo et al., 2024; Tajadura-Jiménez et al., 2015b).
+Last but not least, we recognize that the value of audio personas depend on how widely they are adopted.
+We also acknowledge that the current headphone-based prototype poses a technology barrier, limiting the augmented social experience to users with headphones.
+However, audio personas do not have to be a headphone-delivered experience.
+For example, personal speaker-based (e.g., ”boomboxes,” (Boyer, 2014) popular in the 1980s and 90s) or clothing-integrated audio personas could achieve more universally accessible experiences.
+Individual difference to audio stimuli.
+Analysis of 48 participant-generated designs indicates that audio personas offer rich design possibilities.
+Participants explored a wide range of audio content with varied emotional valence and highlighted diverse patterns for localizations, triggers, and dynamics.
+As designers continue to explore the design space of audio personas, one consideration to be noted is individual differences to audio stimuli.
+For instance, certain sounds might be deemed positive by some people while deemed as negative by others.
+Thus, communication through audio personas may break down if senders and receivers hold different mental models of the same sounds.
+A potential solution is to learn from usage patterns and construct an audio persona library that includes designs with commonly understood meanings.
+This is similar to emoji libraries commonly seen on social media platforms.
+Building on such a shared audio persona library, designers could then offer customization options to support personalized expression.
 6.3. Limitations
@@
 These improvements would enhance the accessibility and practicality of audio persona systems in everyday settings.
-Study #1 interviewed designer-participants to gain initial insights of audio personas usage, following prior work (Liu et al., 2023).
-A larger and more diverse sample, particularly participants without a technical background, is needed to understand the preferences of the broader population.
-Additionally, while Study #1 elicited an initial set of designs, it did not fully explore the challenges of scaling audio personas, such as in large conference settings.
-Future studies should invite designers to investigate techniques like audio mixing for effective deployment in large-scale, multi-user contexts.
-Study #2 employed a controlled, in-lab experiment to investigate how audio personas influence the perception of others, exploring the potential uses identified in Study #1 from the receiver’s perspective.
+Study #1 employed a controlled, in-lab experiment to investigate whether audio personas social perception.
 While this design provided valuable insights into how key design factors of audio personas (e.g., body-anchored, audio valence) influence social perception, it has a few limitations.
@@
 Similarly, researchers should investigate the effects of audio personas without the explicit framing of ”studying how audio impacts social interactions” to encourage naturalistic reactions.
-Second, the study used object-anchored sounds as the control condition, which allows us examine the exact effect of body-anchoring in audio personas.
+Second, the study used object-anchored sounds as the control condition, which allows us to examine the exact effect of body-anchoring in audio personas.
 However, future work should compare audio persona with other baseline, such as no audio or non-spatialized sound.
-Third, in Study #2, we used only nature and animal sounds, but a wider variety of content should be explored, such as human and mechanical sounds.
+Third, in Study #1, we used only nature and animal sounds, but a wider variety of content should be explored, such as human and mechanical sounds.
 This study also only focused on a single style of sound characteristics, investigating variations in localization, triggers, and dynamics of audio personas will provide new insights of how audio personas influence social perception.
@@
 Ultimately, the effects of audio personas in multi-user contexts and real-world social interactions beyond the lab should be explored.
+Study #2 interviewed participants experienced in audio design to gain initial insights of audio personas usage, following prior work (Liu et al., 2023).
+A larger and more diverse sample, particularly participants without a technical background, is needed to understand the preferences of the broader population.
+Additionally, while Study #2 elicited an initial set of designs, it did not fully explore the challenges of scaling audio personas, such as in large conference settings.
+Future studies should invite designers to investigate techniques like audio mixing for effective deployment in large-scale, multi-user contexts.
 6.4. Future Directions
-To deploy audio personas in real-world, multi-user contexts, we lay out several future directions.
-Social identities are multifaceted and adapt fluidly based on contexts (Tajfel, 1974), suggesting that users may wish to have multiple audio personas.
-This opens up opportunities to explore interaction techniques that enable users to seamlessly create, project, and switch between audio personas tailored to different contexts.
-Additionally, as users both project and receive audio personas, the control over auditory experiences becomes essential.
-This includes the ability to choose which audio personas to hear and to start or stop listening as desired.
-Future research could investigate user preferences and interface designs that offer such control, such as opt-in and opt-out mechanisms.
-While audio personas were suggested to be useful in many contexts in Study #1, more investigations are needed to understand situations where audio personas should not be used.
-For instance, when the user needs to focus on something auditory and not be distracted by other signals, audio personas might not be well-suited.
-Additionally, quantitative understanding is needed to assess the effectiveness of audio personas in conveying various types of messages.
-This would help users leverage audio personas to efficiently communicate intended messages and avoid miscommunication that might be introduced by audio-only interactions.
+To further explore the potential of audio personas, future research should investigate their use in more naturalistic settings.
+This would extend the findings from our in-lab study and build toward the potential use cases identified in the design study.
+One promising direction is the deployment of audio persona systems at conferences.
+Conferences are semi-public-private environments where attendees often seek to connect with new people over the course of a few days.
+Such settings offer a unique opportunity to explore the use cases and social functions of audio personas in a scalable yet manageable way.
+Moreover, they provide a valuable platform for studying how individuals manage multiple audio personas to present themselves differently across varying contexts (Tajfel, 1974), as well as for exploring design strategies to support hearing multiple audio personas in shared spaces.
+As audio personas move beyond in-lab environments into more naturalistic settings, another direction for future work lies in exploring the psychological dynamics underlying their use.
+In the in-lab study, we were able to control the behavior and facial expressions of research assistants to remain neutral, isolating the effect of audio personas.
+However, in real-world scenarios, a person’s facial expressions and body movements are dynamic and may not always align with the tone conveyed by their audio persona.
+For example, someone might display a cheerful facial expression while their audio persona sounds sad.
+This kind of mismatch raises important research questions about the underlying processes of social perception when additional—and potentially conflicting—cues are introduced.
 As future work explores more complex audio personas, leveraging knowledge in psychoacoustics offers great potential.
 For example, the best human auditory localization performance occurs within - in the horizontal plane and - in elevation (Carlini et al., 2024).
-In Study #1, we observed that designers localized audio personas to specific body parts to create motion-specific experiences, aligning well with these perceptual thresholds.
+In Study #1, we observed that audio designers localized audio personas to specific body parts to create motion-specific experiences, aligning well with these perceptual thresholds.
 Further research could investigate computational models to support the creation of perceptually distinct audio personas.
 This becomes particularly important to prevent sensory confusion in multi-user contexts.
-Lastly, we recognize that the value of audio personas depend on how widely they are adopted.
-We also acknowledge that the current headphone-based prototype poses a technology barrier, limiting the augmented social experience to users with headphones.
-However, audio personas do not have to be a headphone-delivered experience.
-For example, personal speaker-based (e.g., ”boomboxes,” (Boyer, 2014) popular in the 1980s and 90s) or clothing-integrated audio personas could achieve more universally accessible experiences.
-We see diverse opportunities to deploy audio personas and examine their implications for social interactions.
 7. Conclusion
 We introduced Audio Personas, a novel concept to augment face-to-face interactions with body-anchored audio cues.
-Through a formative study with designers, we uncovered potential use cases and elicited design patterns.
-Our in-lab study further revealed that audio personas can influence how people formed social impressions.
+Our in-lab study revealed that audio personas can influence how people formed social impressions.
 Individuals with a positive audio persona were deemed as more socially attractive, likable, and less threatening than ones with a negative audio persona.
+Through a design study with audio designers, we further uncovered potential use cases and elicited design patterns.
 Audio personas open new possibilities for engaging dynamic multisensory channels to enhance face-to-face social interactions and this paper laid out key building blocks for future explorations.
@@
 We would like to thank Keshav Rastogi, Muhammad Ali, and Brian Beams for their support on the user study logistics.
-We thank Anna Matsumoto and Emmanuel Corona-Moreno for their help on the video production.
-We would also like to thank everyone who provided feedback on the paper draft: Danyang Fan, Han Guo, Eugy Han, Saehui Hwang, Monique Santoso, Tara Srirangarajan, Olivia Tomassetti, Ahad Rauf, Savannah Cofer, Portia Wang, Yifei Cheng, Yiran Zhao, and Ran Zhou.
+We thank Anna Matsumoto and Emmanuel Corona-Moreno for their help on the paper video.
+We would also like to thank everyone who provided feedback on this project: Ethan Buck, Prof.
+Ana Tajadura-Jiménez, Danyang Fan, Han Guo, Eugy Han, Saehui Hwang, Monique Santoso, Tara Srirangarajan, Olivia Tomassetti, Ahad Rauf, Savannah Cofer, Portia Wang, Yifei Cheng, Yiran Zhao, and Ran Zhou.
 The lead author of this work is supported by the Knight-Hennessy Scholarship.References
@@
 - Fiore (1993) Ann Marie Fiore. 1993. Multisensory integration of visual, tactile, and olfactory aesthetic cues of appearance. Clothing and Textiles Research Journal 11, 2 (1993), 45–52.
+- Frid et al.
+(2016) Emma Frid, Roberto Bresin, Paolo Alborno, and Ludvig Elblaus.
+2016.
+Interactive sonification of spontaneous movement of children—cross-modal mapping and the perception of body movement qualities through sound.
+Frontiers in neuroscience 10 (2016), 521.
 - Fuller et al.
@@
 https://doi.org/10.1145/3170427.3188589
-- Goffman (2023) Erving Goffman. 2023. The presentation of self in everyday life. In Social theory re-wired. Routledge, 450–459.
 - Goffman et al. (2002) Erving Goffman et al. 2002. The presentation of self in everyday life. 1959. Garden City, NY 259 (2002).
@@
 113–121.
+- Hall (1963) Edward T Hall. 1963. A system for the notation of proxemic behavior. American anthropologist 65, 5 (1963), 1003–1026.
 - Hall (1966) Edward T Hall. 1966. The hidden dimension. Garden City (1966).
@@
 Psychology of Music 43, 1 (2015), 75–85.
+- Hayduk (1983) Leslie A Hayduk. 1983. Personal space: Where we now stand. Psychological bulletin 94, 2 (1983), 293.
 - Hong et al.
@@
 1–20.
+- Jung (2014) Carl Gustav Jung. 2014. Two essays on analytical psychology. Routledge.
 - Juslin and Västfjäll (2008) Patrik N Juslin and Daniel Västfjäll. 2008. Emotional responses to music: The need to consider underlying mechanisms. Behavioral and brain sciences 31, 5 (2008), 559–575.
@@
 Interacting with computers 22, 5 (2010), 336–343.
-- Oxford English Dictionary (2023) Oxford English Dictionary. 2023. persona (n.), sense 2.a. https://doi.org/10.1093/OED/1145536616
 - Pijanowski et al.
@@
 ACM Transactions on Multimedia Computing, Communications, and Applications (TOMM) 14, 3 (2018), 1–22.
+- Singh et al.
+(2024) Aneesha Singh, Marusa Hrobat, Suxin Gui, Nadia Bianchi-Berthouze, Judith Ley-Flores, Frederic Bevilacqua, Joaquin R Diaz Duran, Elena MÁrquez Segura, and Ana Tajadura-Jiménez.
+2024.
+Pushed by Sound: Effects of Sound and Movement Direction on Body Perception, Experience Quality, and Exercise Support.
+ACM Transactions on Computer-Human Interaction 31, 4 (2024), 1–36.
 - Spence (2013) Charles Spence.